### v1

# Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis

Abstract
Clinical case reports and discharge summaries may be the most complete and accurate summarization of patient encounters, yet they are finalized, i.e., timestamped after the encounter. Complementary data structured streams become available sooner but suffer from incompleteness. To train models and algorithms on more complete and temporally fine-grained data, we construct a pipeline to phenotype, extract, and annotate time-localized findings within case reports using large language models. We apply our pipeline to generate an open-access textual time series corpus for Sepsis-3 comprising 2,139 case reports from the Pubmed-Open Access (PMOA) Subset. To validate our system, we apply it on PMOA and timeline annotations from I2B2/MIMIC-IV and compare the results to physician-expert annotations. We show high recovery rates of clinical findings (event match rates: O1-preview–0.755, Llama 3.3 70B Instruct–0.753) and strong temporal ordering (concordance: O1-preview–0.932, Llama 3.3 70B Instruct–0.932). Our work characterizes the ability of LLMs to time-localize clinical findings in text, illustrating the limitations of LLM use for temporal reconstruction and providing several potential avenues of improvement via multimodal integration.
1 Introduction
The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria (Kyriazopoulou et al., 2021) and phenotyping studies (Seymour et al., 2019). In the critical care literature, sepsis is seen as a heterogeneous disease, a pathway arrived upon from a multitude of infectious origins, anatomical sites, and comorbid profiles. There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems (Henry et al., 2022), variation in prediction utility (Kamran et al., 2024), and endotype characterization (Noroozizadeh et al., 2023).
Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis (Moldwin et al., 2021). A more complete alternative is the discharge summary, or more generally, the case report, in which a clinician documents the clinical findings pertinent to the case and care management. The impracticality of using the case report is that it is only written in full after the encounter is over, i.e., with a less granular timestamp than the events contained within. In principle, having a more complete, fine-grained temporal record could enable better temporal analyses of patients with sepsis, leading to better understanding and treatment improvements.
At present, the (semi-)public critical care repositories at the research frontier contain high resolution structured signals and low resolution textual signals. While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts (Sun et al., 2013) rather than on event timing. Those that have focused on timing (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024) used excerpted summaries and relatively small sample sizes.
In order to train sepsis models that can consider sources of information references in free-text and that have fine time-granularity, we seek to generate larger samples of textual time series. To do so, we introduce a novel Sepsis-3 textual time series corpus from the Pubmed Open Access (PMOA) Subset. We adopt the LLM-as-annotator approach, evaluating its ability to extract clinical findings and its ability to accurately associate them with timestamps. We augment existing clinical annotations with our own on PMOA, and contrast LLM annotations with our own and that of Frattallone et al. (2024). To provide checks on their performance at scale, we consider a larger subsample for inter-LLM annotator agreement, and show the performance levels across several model sizes, which provides valuable context for cost-based analyses. Finally, these assessments provide evidence of the quality of the LLM annotations and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
1.1 Related work
Numerous critical care and sepsis datasets are available (semi-)publicly, e.g., Thoral et al. (2021); Pollard et al. (2018), including several with de-identified clinical discharge summaries: MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023). Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data (Johnson et al., 2018), although structured data is often highly missing, leading to suboptimal phenotypes (Seinen et al., 2025). Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models (Kamran et al., 2024).
Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024). In the case of Leeuwenberg and Moens (2020), the event spans are already given, and focus on an excerpt of the full discharge summary, so we focus our direct comparison on that of Frattallone-Llado et al. (2024). In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept (Uzuner et al., 2011), which allows for greater specificity of the event in the absence of additional context. Additionally, as compared to Frattallone-Llado et al. (2024), we focus on using text alone for clinical time series construction, which is necessary for the Pubmed Open-Access Subset since no complementary data source is available.
Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization (Van Veen et al., 2024). In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models (Jeong et al., 2024). In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to the performance of the model.
1.2 Contributions
Our work highlights the following contributions: (1) we present the first textual time series corpus for sepsis, comprising 2,139 open-access case reports identified as having sepsis; (2) we develop an analytic pipeline to identify Sepsis-3 from clinical text and to extract time-associated clinical findings; (3) we annotate time-associated clinical findings on a random subset of the case reports; (4) we develop and adapt evaluation criteria for textual time series, where matched clinical findings are not expected to be identical text strings; and, (5) in multiple datasets, we validate the performance of our pipeline, showing strong annotation performance compared to manual annotation and improved temporal results compared to existing methods.
1.3 Generalizable Insights
Access to the best temporal information available is critical in sepsis prediction tasks and has major implications in clinical decision making and early warning. The analysis pipeline we validate and corpus we introduce enable the use of textual data with temporal granularity, expanding the available features for use in a multitude of tasks. For example, the recent investigation in JAMA on vasopressin treatment policies using offline reinforcement learning uses hourly structured data in its learning, its reweighting analyses, and its validations (Kalimouttou et al., 2025). Access to textual features would augment the information that could affect its predictions, recommendations, and its adjustment for measured confounding, which may alter and refine its main findings. By adopting or extending our approach, the temporal resolution of the clinical event findings available only in text could be used and accounted for. As observed in Seinen et al. (2025), this may be a large fraction of concepts, with only 13% of textual events identified in structured form in their study.
2 Methods
We define the task specification, relevant terms, the extraction and annotation procedure, and the assessments.
Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3. We define the term textual time series as a list of clinical findings each with an associated timestamp (which may be absolute or relative to time of case presentation) pertaining to an individual. A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
i2m4
[…]
Yes 2 0
No 0 18
L3.1 8B Yes No
Yes 2 7
No 0 11
sepsis-100
[…]
Yes 78 7
No 8 7
L3.1 8B Yes No
Yes 81 12
No 5 2
Figure 1: Flowchart of the PMOA T2S2 corpus construction pipeline (left). Confusion matrices (right) for Sepsis-3 phenotype in i2m4 (top) and sepsis-100 (bottom) datasets, with preferred annotator in column position.
Compared to clinical concept annotations (Uzuner et al., 2011; Sun et al., 2013), we choose to expand the text span determination to increase the specificity of the meaning attached to the finding. We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications (Uzuner et al., 2011):
  1. 1.
    Clinical findings are not limited to one prepositional phrase following a markable concept (e.g., “pain in chest that radiates substernally”). Instead of splitting into concepts “pain in chest” and possibly “substernal radiation”, where the former lacks specificity and the latter lacks origin, a single finding is preferred.
  2. 2.
    Related conjunctions should be split into component findings. Instead of “metastases in the liver and pancreas”, we prefer two more specific clinical findings “metastasis in the liver” and “metastasis in the pancreas”.
The other design choice we make is to alter the time expression to disambiguate timing by capturing the start time of the perception of the clinical event. For example, if a patient has a “three-day history of fever”, the clinical event “fever” is assigned to -72 hours, since it is first perceived then. If the patient has a “history of smoking”, the start time is fully ambiguous, so the time of perception is the time that smoking is assessed, e.g., at the start of the encounter, and the event is “history of smoking”, rather than “smoking”. To restate, our clinical finding is defined as a contiguous text span from the text corresponding to a temporally localizable event pertaining to the patient, modified for conjunction splitting and temporal perception as above.
Given this specification, our subtasks were: case report detection, Sepsis-3 detection, clinical finding extraction, and timestamp determination. We constructed a pipeline for these steps and their assessment (Figure 1 (left), showing the process for the PMOA dataset, with an analogous process used for i2m4).
2.1 Data
We used two data sources: (1) the absolute timeline annotations from Frattallone-Llado et al. (2024), and (2) the Pubmed Open Access subset (PMC Open Access Subset, 2024) (see Table 1).
2.1.1 i2m4 Dataset
The absolute timeline annotations (Frattallone-Llado et al., 2024) came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts. For our analysis, we used the lower bound of the interval (specified as a mean time) as it denotes the time when the finding may first be observed. We used only the probabilistic annotations, since the bounded annotation lower bound could precede the actual occurrence of the finding. Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated in the same way)—hence we call this dataset “i2m4”.
Table 1: Dataset summary. Bold cells indicate contributed labeling
Corpus i2m4 Sepsis-10   Sepsis-100
N (unlabeled) 310; 331,794 2,139 2,139
N (labeled) 20 10 100
Source i2b2; MIMIC IV PMOA PMOA
Note type discharge summary case report case report
Labeler clinicians clinicians O1-preview
Phenotyper clinicians clinicians QwQ-32B-Preview
Methods O1-preview O1-preview
DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1
L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct
L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct
Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B
 Instruct v0.1  Instruct v0.1  Instruct v0.1
L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct
Ablations No role No role No role
Base: No conjunction No conjunction No conjunction
L3.3 70B Instruct  instruction  instruction  instruction
No few-shot prompting No few-shot prompting No few-shot prompting
Sensitivity analyses With upper bound With upper bound With upper bound
With upper bound + With upper bound + With upper bound +
 i2b2 event typing  i2b2 event typing  i2b2 event typing
For Sepsis-3 determination, we had a clinician review the discharge summaries and assign each case as meeting the Sepsis-3 definition. To develop a computational approach, we constructed a prompt specific to Sepsis-3, which describes the qSOFA procedure for determining a Sepsis-3 diagnosis, and used it as query to Llama-3.1-8B-Instruct (L3.1 8B) and QwQ-32B-Preview (Qwen). We used the phenotyping prompt:
[…]
We constructed confusion matrices to contrast the Sepsis-3 phenotyping abilities of the LLM-query against clinician determinations (Figure 1, top right).
2.1.2 PMOA (sepsis-10, sepsis-100) Dataset
The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications (PMC Open Access Subset, 2024). Because the majority of these publications are not case reports, we built an extraction pipeline to recover critical care case reports, i.e., a clinical research informatics critical care cohort.
We used the PMOA Subset (version as of April 25, 2024). We removed the header and footer of the free-text which contained the Title, Abstract, and References sections, and retained the ==== Body section. To identify potential case reports, we required that the free-text match the case-insensitive regular expressions case (report|presenta) and year-? ?old. We contrasted the output of the query with PubMed metadata for case reports and found our regular expression approach had substantially higher specificity during inspection of a sample of 20 random reports.
To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||||septic) and (critical||||intensive) care. Then, we used the an LLM query to extract sepsis cases, as well as the number of cases present, and the age and gender. We used GPT-4o and Llama-3.1-8B-Instruct for the phenotyping query screen. Reports were omitted if the number of cases identified was not equal to 1. As a screen, we chose to include the case report if either GPT-4o or Llama-3.1-8B-Instruct identified the report as having Sepsis-3.
This process resulted in 2,139 identified Sepsis-3 case reports, respectively, among which 10 and 100 were sampled for further review. The sepsis-10 sample was annotated by a clinical expert to provide a “ground truth” labeling, and the sepsis-100 sample was used as a larger sample for testing. In sepsis-100, to better assess Sepsis-3 presence, we used the QwQ-32B-Preview model and Sepsis-3 phenotyping prompt above (since it had strong performance characteristics in the i2m4 dataset) to validate the GPT-4o and Llama-3.1-8B-Instruct responses (Figure 1, bottom right).
2.2 Textual Time Series Annotations
For our three datasets—i2m4, sepsis-10, and sepsis-100—we used the following LLMs: O1-preview (o1-preview-2024-09-12), DeepSeek-R1-UD-IQ1_S, Llama-3.3-70B-Instruct, Llama-3.1-70B-Instruct, Mixtral-8x7B-Instruct v0.1, and Llama-3.1-8B-Instruct. For data privacy reasons, we did not submit any i2m4 data queries to O1-preview and ran the other models locally. For the sepsis-100 dataset, we did not have access to clinician annotations, so our inter-LLM annotator assessments used the O1-preview annotations as reference. We attempted to use QwQ-32B-Preview and Qwen-2.5-72B-Instruct for clinical finding annotations but were unable to reliably extract query results. We adapted the Llama query using the system/user/assistant query template. The main annotation prompt used is the following:
You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamp are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report.
An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment. So the event of fever and rash happen four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672. diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the time stamp is 0 hours, since it happens right at admission. DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0. Then the output should look like
18 years old |||| 0
male |||| 0
admitted to the hospital |||| 0
fever |||| -72
rash |||| -72
acne |||| -672
minocycline |||| -672
increased WBC count |||| 0
eosinophilia|||| 0
systemic involvement|||| 0
diffuse erythematous or maculopapular eruption|||| 0
pruritis |||| 0
DRESS syndrome |||| 0
fever persisted |||| 0
rash persisted |||| 0
discharged |||| 24
[…]
For sensitivity analysis to the prompt used, we omitted the color-coded parts in turn to give us the following ablations: no-role playing (no-role), zero-shot prompting (zero-shot), and no conjunction expansion (no expansion). We also tested with augmentations, where we add to the query a request for the time interval (Interval) or time interval and i2b2 event type (Interval+Type), where event type is one of: Factual, Possible, Hypothetical, Conditional, Negated, Historical, Uncertain. For the augmentations, both the original prompt’s instructions and few-shot examples were modified accordingly.
For the i2m4 dataset, we also compared against a BERT-based baseline (BERT 5-CV), our adaption of the Frattallone et al. model (2024). In their work, they transformed the temporal prediction task into multiclass classification and used a multilayer perceptron (MLP) head with a softmax and cross-entropy loss for training. To predict relative time (hours from case presentation), we made the following adaptations. Similarly to the classifier module, we used a 3-layer MLP regressor head from the output of the BERT event span features and output the prediction means and standard deviations. The regression task (predicting the lower bound) was optimized using L1 loss. We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span (Zhong and Chen, 2021) and removed the dropout layers inside the MLP heads, leaving only one at the final layer. Regarding the timestamps used for prediction, while many events occurred near the admission time, a subset extended months or years away, complicating the design of a regressor capable of both spanning wide magnitudes and maintaining precision for the majority of the means close to the admission time. To address this, we filtered the findings by their magnitudes, retaining them if the means were within two weeks of the admission time and the standard deviations were under two weeks. Filtering reduced the dataset size (in number of findings) from 3,472 to 3,350. Training was carried out for 50 epochs with batch size 32 and learning rate 3e-5. The model training used 5-fold cross-validation.
2.3 Assessments
We assessed the textual time series with the following methods: event match rate, concordance (c-index), median absolute time discrepancy, and log-time discrepancy. To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix B for pseudocode). Specifically, the recursive best match was used to reduce erroneous surjective matches, i.e., multiple events from the predicted event list mapping to the same event in the reference list. For each reference event, if multiple predicted events matched, the lowest distance match was selected (with event ordering used to break distance ties), and both the predicted and referent events were removed from the lists for the recursive call to match events. Then all matched events from the recursive calls were returned. To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) (Deka et al., 2022). Upon manual review of event matches, we selected option 3 with a cosine similarity cutoff of 0.1.
Among event matches, we then considered the time-ordering and time-discrepancy between label timestamps and predicted timestamps. We computed the concordance, or c-index, which corresponds to the probability of correctly time-ordering a random pair of predicted findings. The median of absolute error of time discrepancies was used to avoid the outsize impact of large discrepancies that would be captured in the mean absolute error. For computing log-time discrepancies, we tranformed the non-negative discrepancies with the function: $\log(1+\Delta t)$, where $\Delta t$ is the time discrepancy (absolute) per clinical finding in hours. Finding-level cumulative time discrepancy plots order the discrepancies from small to large to show the fraction of discrepancies being smaller than some discrepancy threshold. Then, the area under the log-time curve (AULTC) is the area under the log-time curve of the cumulative time discrepancy plot, which we describe next.
2.3.1 Log-Time Cumulative Distribution Function
Given we have $k$ event matches combined across case reports, we want to compute a measure of the discrepancy between the predicted times $t^{p}_{i}$ and the manual reference times $t^{r}_{i}$ for $i={1,\dots,k}$ that captures the quality of time recovery across time scales.
We define the log-time cumulative distribution function as follows. For each matched event $i$, compute the discrepancy $s_{i}=|t^{p}_{i}-t^{r}_{i}|$. then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours):
$s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$
Apply the log transformation:
$x_{i}=\log(1+s^{\prime}_{i})$
Sort $x_{i}$ in ascending order:
$x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$
The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$:
$F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$
where $\mathbf{1}$ is the indicator function.
Then, we get the following boundary values: $F(0)\geq 0$ (depending on the smallest discrepancy) and $F(\log(1+S_{\text{max}}))=1$.
With the the CDF defined, we compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
AULTC $=\frac{1}{\log(1+S_{\text{max}})}\Bigg{[}\Big{(}\sum_{i=1}^{k}(x_{(i)}-x_{(i-1)})\frac{i}{k}\Big{)}+(\log(1+S_{\text{max}})-x_{(k)})\times 1\Bigg{]}$
where $x_{(0)}=0$. With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area.
We provide additional remarks regarding the need for explicit bound specification of the log-time CDF and AULTC and the non-convexity of the average log discrepancy in Appendix A.
2.4 Annotation Quality Review
To add to the quantitative assessment of annotations at the event-time level, we also verified the annotation quality at the case-report level in our annotation quality review. To inspect the quality of the annotations, we had our physician annotator review the annotations of the Sepsis-10 dataset that the LLM models produced. The annotator was asked to rank-order the annotations (as well as the manual annotation file), with rank 1 being the best, and rank 7 being the worst. The annotator also placed the annotations into quality levels of Excellent, Good, Acceptable, and Poor. The guideline for placing the annotations into each quality level was:
Excellent: High quality extraction. May contain up to 1 clear error in extraction or timing; extracts all remaining important findings; ordering and timing are plausible.
Good: Extracts the important findings, almost always gets the ordering correct, and greater than three-quarters of timings are plausible.
Acceptable: Extracts the important findings, mostly gets the ordering correct, timing is reasonable for more than half of the timing groups.
Poor: Fails to meet criteria for Acceptable.
Annotations of the same quality level were assigned their lower rank (closer to 1). After placing in approximate relative rank, the annotator was asked to pairwise compare all adjacent annotations to ensure relative rankings and ties. The annotator was asked to document pertinent findings and errors in free text.
Table 2: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets.
i2m4 sepsis-10 sepsis-100
Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
O1-preview 0.76 0.932 22 0.796
DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801
L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790
Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737
L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705
BERT 5-CV 0.704 34 0.520
3 Results
Our analysis resulted in T2S2, a new open-access textual time series corpus for Sepsis-3, comprising 2,139 case reports. Demographically, the corpus is 58% male, 42% female, with ages ranging from 0 to 111 and a mean age of 49 (IQR: 32-65). The sepsis-10 manual annotations are of 10 files: PMC3075162, PMC4778217, PMC6238558, PMC7576661, PMC9136437, PMC9355071, PMC9552858, PMC10370906, PMC10556716, and PMC10629858, resulting in 801 reference events. An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix C.
3.1 Phenotyping Results
For phenotyping the i2m4 dataset for sepsis, we found just two case reports of patients with sepsis. The QwQ-32B-Preview response had perfect accuracy (100%), whereas the Llama-3.1-8B-Instruct accuracy was 13/20 (65%). For the sepsis-100 dataset, where QwQ-32B-Preview outputs were labels, the relatively faster models GPT-4o and Llama-3.1-8B-Instruct agreed 84% and 83% of the time, respectively (Figure 1, bottom right). Although the QwQ-32B-Preview responses were repetitive and verbose when making a Sepsis-3 diagnosis, they were visibly based on external foundation knowledge not provided in the prompt, e.g., through use of SOFA rather than qSOFA for phenotyping.
3.2 Annotation Results
The results in Table 2 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets. In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list (Frattallone-Llado et al., 2024)), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding. Among the matches, Llama 3.3 and DeepSeek R1 performed best for time predictions (trading off event matching with temporal performance), and substantially outperforming the baseline BERT model, particularly with respect to time discrepancy. However, the overall results on i2m4 were considerably lower than in the PMOA datasets, including in terms of time ordering and time discrepancy. In contrast to concordances of 0.707 in i2m4, in sepsis-10, O1-preview and Llama 3.3 were excellent time-ordering annotators, both with concordances of 0.932.
Table 3: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for i2m4, sepsis-10, and sepsis-100.
i2m4 sepsis-10 sepsis-100
Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
Ablations
No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789
No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793
Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795
Augments
Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791
Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
To examine the results in greater detail, we show figures of the cumulative distribution function of event matches and concordance boxplots (Figure 2) for the sepsis-10 dataset. When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event that the prediction method did not. The event match rates in Table 2 correspond to the match rate at the cosine distance threshold in Figure 2. One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positive matches, which degrade the temporal performance characteristics. The concordance box plots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower, which is a consistent finding throughout our analysis.
To better visualize the time discrepancies, we plot cumulative time discrepancy plots (Figure 2) that show, for example, that 44% of the O1-preview annotations are exact time matches and that 80% are within 1 week of the label timestamp. Breaking it down further (Figure 2, right), we can see that among events occurring within 1 day of presentation (but more than 1 hour), 90% of O1-preview annotations are within a 1-week time error. This drops to 60% when considering events up to a year from presentation. Problematically, for labeled events more than 1 year from presentation, only O1-preview and DeepSeek R1 ever identify the label time to within 1 year exclusive (Figure 2, right).
The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3). Our original prompt appears to be the strong performer overall, showcasing the value in particular of few-shot prompting, but also the value of the technical specification of how to treat conjunctions and the modest benefit of role-playing. Surprisingly, when comparing the main L3.3 70B model responses to the augmentations responses, it appears that requesting timestamp intervals (Interval) modestly increases the event match rate. However, the temporal performance characteristic degrades, suggesting a possible request overload on the model or an increase in false positive event matches. The performance of the typing method (Interval + Type) is further degraded. Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix D. To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix E. This analysis provides additional insight into the factors contributing to variations in model performance for event match rate and c-index across datasets.
Figure 2: Event match cumulative distribution function (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
3.3 Annotation Quality Results
Inspecting the clinical quality of the annotations, we report the mean rank, top-1 fraction, top-1-LLM fraction, and quality level in Table 4. There is a clear performance ordering, with O1-preview annotations being a close second to the manual annotations, followed by DeepSeek R1 IQ1 and Llama 3.3. While the Llama 3.3 annotations are found to be acceptable a high percentage of the time (90%), the other Llama and Mistral approaches fail to achieve Acceptable performance in excess of 50%. Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2 and Figure 2 (left and middle right)). Our clinical reviewer commented that the Llama 3.3 discrepancies for large time horizons, i.e., years, were large, which is corroborated in Figure 2 (right, bottom subpanels).
The annotation review feedback noted a wide variety of errors; here we present several errors modes identified during quality review.
Events with duration. Example: “10-kg weight loss” from “he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss.” This example from case report PMC10629858 references a gradual event that has been ongoing for 2 months. By the annotation guideline for selecting the start time of events with duration, the time -168 (what O1-preview annotates) is correct. However, the event can also be considered as “history of 10-kg weight loss” via the “history of” guideline with time 0 (what L3.3 70B annotates). Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3 and Figure D.1).
Events requiring temporal reasoning or composition. Example: “passed away” from “he was re-admitted to the medical ICU for severe sepsis and multiorgan failure and passed away around 6 months after his initial diagnosis with NHL, despite maintaining a remission status.” This requires a two-step inference, (1) timing of NHL diagnosis, and (2) adding “around 6 months” to that time. While this is a simple example, more complicated chains of relative timings were noted, and possibly relative to other events whose timings are also uncertain. As black-box methods, there is no explicit probabilistic specification for these chains, and empirically we see variable responses. For this case, the manual annotation is: 4383 (about 6 months), O1-preview gives 4368 (about 6 months), DeepSeek R1 gives 4032 (about 6 months), and L3.3 70B gives: 1440 (about 2 months; compared to other events the temporal ordering is correct, but the relative timing is off by months).
4 Discussion and Conclusions
We developed an analytical pipeline for processing free text biomedical corpora to create textual time series for Sepsis-3 using LLM and showed strong performance characteristics, including good match rates, high concordance and low time discrepancies. We plan to release the full cohort annotations upon acceptance, providing a novel Sepsis-3 cohort with rich textual information, annotated with timestamps, which may be useful for downstream tasks such as risk forecasting, causal reasoning, and disease trajectory characterization. In PMOA samples sepsis-10 and sepsis-100, performance levels were strong with respect to clinician annotations (sepsis-10) and inter-LLM agreement (sepsis-100).
Table 4: Sepsis-10 annotation quality review
Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable
Manual 1.2 0.9 1 1 1
O1-preview 1.5 0.6 0.9 0.9 1 1
DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8
L3.3 70B 3.8 0 0 0 0.2 0.9
L3.1 70B 4.8 0 0 0 0.2 0.5
Mixtral 8x7B 5.4 0 0 0 0.1 0.2
L3.1 8B 6.5 0 0 0 0 0
There are several limitations to our work. First, our product, the textual time series corpus for sepsis, is a clinical research informatics cohort, and possesses its characteristic limitations. It should not be considered a representative sample, since case report corpora possess biases, including but not limited to publication biases, rare finding over-representation, and revisionism. Second, while we analyzed multiple data sources, we are limited by the number of clinician annotations that are challenging and expensive to obtain. Third, the substantial use of LLMs in the analytic pipeline can make error analysis more difficult, as their errors can be subtle and inaccurate.
Regarding generalizability to other medical conditions besides sepsis, our methodology could be extended in a straightforward manner. Our motivation for choosing sepsis as the disease focus was to expand the analysis from the i2m4 dataset to a similar population in the PMOA corpus and maintain a well-defined phenotype. Sepsis is a sensible choice as it is one of the most prevalent and important diagnoses in critical care medicine. However, there are several concerns when generalizing to other conditions. For one, most other conditions are less acute in nature, meaning that the time distributions may be wider. Also, the amount and type of information in such case reports may vary. Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when you mix intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales will not dominate the performance measure (because of the time discrepancy log scaling). Another generalizability concern is that for some conditions, less temporal information will be present. This appears to be more acute in imaging studies, where the focus of the case is on the image content at a single snapshot in time. For these types of studies, the longitudinal annotations become less meaningful compared to the image’s information content.
Future work could further characterize the findings, with additional or improved augmentations to the current system. Although our augmentation strategies did not improve overall system performance (Table 3), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation. Our hope is that this corpus, or a future version of it, will be used for downstream applications in sepsis, such as causal modeling, forecasting, and multimodal alignment. As a community, we must use the best data available regardless of its format, and for temporal tasks, this temporal, textual approach offers a complementary corpus to those in frequent use. Further developments in temporal extraction could improve upon our pipeline and could be tailored to one or more of the applications above, each of which place different emphasis on what constitutes a high-quality clinical time series. For example, time-ordering is a necessary condition for causality, whereas the precise timestamp may be of high importance in forecasting and alignment. Finally, while our clinical finding and timestamp pairs attempt to capture the true occurrence and timing of each event, another approach focusing on the timing of perceived occurrences could better uncover how actors process and affect their environments, leading to better characterization for decision analysis, multi-agent systems, and resource allocation. Multimodal integrations could help account for perception differences and design collection strategies to assist early detection of sepsis and sepsis progression.
5 Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster.
Appendix A Log-Time Cumulative Distribution Function
Recall the log-time cumulative distribution function is given as follows:
$F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$
where $\mathbf{1}$ is the indicator function.
We compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
AULTC $=\frac{1}{\log(1+S_{\text{max}})}\Bigg{[}\sum_{i=1}^{k}(x_{(i)}-x_{(i-1)})\frac{i}{k}+(\log(1+S_{\text{max}})-x_{(k)})\times 1\Bigg{]}$
where $x_{(0)}=0$. With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area. The code for the AULTC calculation is given in the function ecdf_auc2(…) in the Supplementary Materials.
Remark A.1.
The time unit and cutoff $S_{\text{max}}$ affect the AULTC calculation, so they must be specified when reporting the AULTC.
Because of the $\log(1+\cdot)$ transformation (rather than the $\log(\cdot)$ transformation which is undefined for zero time-error discrepancy), the discrepancies are non-linearly shifted in the log scale. The non-linear shift adjusts the relative widths of the step function (particularly for small discrepancies) changing the area calculation. The cutoff $S_{\text{max}}$ affects the normalization factor. Therefore, these values should be chosen based on practicalities for the application. In our case, we chose hours at the time unit because the only sub-hour descriptions reported were several Apgar scores at the minute-level.
Remark A.2.
The average log-time discrepancy is non-convex.
This can be seen from the observation that the $\log(1+\cdot)$ function is “spikier” than the L1 function. More formally, the log-time discrepancy has a higher curvature around zero error compared to the L1 loss. Consider the second derivatives of the loss functions with respect to the discrepancies $s=|t^{p}-t^{r}|$. For L1 loss $L_{1}(s)=s$, the second derivative $\frac{d^{2}L_{1}}{ds^{2}}=0$ for $s>0$. For the log-time loss, $L_{log}(s)=\log(1+s)$, the second derivative is $\frac{d^{2}L_{log}}{ds^{2}}=-\frac{1}{(1+s)^{2}}$. At $s=0$, the second derivative is $-1$, while the second derivative of the L1 loss is 00.
To provide a simple counterexample, define the average log-time discrepancy is given by:
$L(t^{p})=\frac{1}{k}\sum_{i=1}^{k}\log(1+\min(|t^{p}_{i}-t^{r}_{i}|,S_{\text{max}}))$
To prove that this function is non-convex, we need to show that there exist $t^{p1},t^{p2}$ and $\lambda\in(0,1)$ such that
$L(\lambda t^{p1}+(1-\lambda)t^{p2})>\lambda L(t^{p1})+(1-\lambda)L(t^{p2})$
Consider the case where $k=1$, $t^{r}=0$, and $S_{\text{max}}=2$. The loss function is $L(t^{p})=\log(1+\min(|t^{p}|,2))$. Let $t^{p1}=0$ and $t^{p2}=3$. Choose $\lambda=0.5$. Then $\lambda t^{p1}+(1-\lambda)t^{p2}=0.5\times 0+0.5\times 3=1.5$. Now we evaluate the loss function at these points:
$L(t^{p1})$ $=L(0)=\log(1+\min(|0|,2))=0$
$L(t^{p2})$ $=L(3)=\log(1+\min(|3|,2))=$
$\qquad\log(3)$
$L(\lambda t^{p1}+(1-\lambda)t^{p2})$ $=L(1.5)=\log(2.5)$
Now we check the convexity condition:
$L(1.5)\leq 0.5L(0)+0.5L(3)$
$\log(2.5)\leq 0.5\times 0+0.5\times\log(3)$
$\log(2.5)\leq\log(\sqrt{3})$
Since $2.5>\sqrt{3}\approx 1.732$, we have $\log(2.5)>\log(\sqrt{3})$. Therefore,
$L(\lambda t^{p1}+(1-\lambda)t^{p2})>\lambda L(t^{p1})+(1-\lambda)L(t^{p2})$
This proves that the average log-time discrepancy is non-convex. This remark illustrates the loss form, in case it is considered for optimization/model training rather than for assessment purposes.
Appendix B Recursive Best Match Procedure
We provide pseudocode for the best match procedure between two lists of strings (Algorithm B.1). For each list, we use the text order, i.e., the order of the events in the annotation files, to break embedding distance ties. We use the cosine similarity for the distance calculation using sentence transformer embeddings from S-PubMedBert-MS-MARCO.
Algorithm B.1 Recursive Best Match
Input :
Two lists: ref (reference events) and pred (predicted events)
Output :
List of best-matching event pairs
MatchEvents(ref, pred)
if ref is empty or pred is empty then
       return []
end if
[…]
       foreach $p$ in pred do
[…]
             end if
[…]
                   end if
[…]
                   end if
             end if
       end foreach
end foreach
[…]
Appendix C Example Annotation
In this section we present an excerpt from PMC10629858, the most recently published case in the sepsis-10 dataset (Abu-Tineh et al., 2023). To illustrate the extraction task and performance characteristics of the LLM annotators, we contrast the manual annotations with that of O1-preview, Llama 3.3, and Llama 3.1 8B.
Excerpt from Abu-Tineh et al. (2023):
A 57-year-old man recently diagnosed with lepromatous leprosy was confirmed with skin biopsy and had been on treatment (rifampicin/clofazimine/dapsone) for 2 months before admission; he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss. On examination, the patient was vitally stable. He had evidence of peripheral lymphadenopathy with a distended abdomen and a positive shifting dullness. A computed tomography scan of his abdomen showed mural thickening of the terminal ileum with significantly enlarged mesenteric lymph nodes, mesenteric fat stranding, and intra-abdominal free fluid, suggesting abdominal granulomatous infection or neoplastic process. …
The patient was planned for consolidation by autologous bone marrow transplant. Unfortunately, with the recurrent bacteremia and sepsis that accompanied the patient’s course due to his immunocompromised state, he was re-admitted to the medical ICU for severe sepsis and multiorgan failure and passed away around 6 months after his initial diagnosis with NHL, despite maintaining a remission status.
Manual annotation
-year-old 0 man 0 diagnosed with lepromatous leprosy -1461 on treatment -1461 rifampicin -1461 clofazimine -1461 dapsone -1461 presented to the hospital 0 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 vitally stable 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan of his abdomen 0 mural thickening of the terminal ileum 0 significantly enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 re-admitted to the medical ICU 4383 severe sepsis 4383 multiorgan failure 4383 passed away 4383
O1-preview annotation
Recently diagnosed with lepromatous leprosy -1440 Skin biopsy confirmed lepromatous leprosy -1440 Started on rifampicin -1440 Started on clofazimine -1440 Started on dapsone -1440 -kg weight loss -168 -year-old man 0 Presented to the hospital 0 Complaints of abdominal distension 0 Complaints of constipation 0 Complaints of vomiting 0 Patient was vitally stable 0 Evidence of peripheral lymphadenopathy 0 Distended abdomen 0 Positive shifting dullness 0 Computed tomography scan of abdomen 0 Mural thickening of terminal ileum 0 Significantly enlarged mesenteric lymph nodes 0 Mesenteric fat stranding 0 Intra-abdominal free fluid 0 Planned for consolidation by autologous bone marrow transplant 2136 Responded well to chemotherapy 2280 Attained complete metabolic remission by positron emission tomography/computed tomography 2280 Re-admitted to medical ICU for severe sepsis and multiorgan failure 4360 Passed away around 6 months after his initial diagnosis with NHL 4368
DeepSeek-R1-IQ1 annotation
years old 0 male 0 lepromatous leprosy -1464 skin biopsy -1464 rifampicin -1464 clofazimine -1464 dapsone -1464 admitted to the hospital 0 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 vital stability 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan 0 mural thickening of the terminal ileum 0 enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 abdominal granulomatous infection or neoplastic process 0 plan for autologous bone marrow transplant 0 recurrent bacteremia 0 sepsis 0 readmission to medical ICU 0 severe sepsis 24 multiorgan failure 24 passed away 4032
Llama-3.3-70B-Instruct annotation
years old 0 male 0 lepromatous leprosy -720 treatment with rifampicin/clofazimine/dapsone -720 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 mural thickening of the terminal ileum 0 enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 abdominal granulomatous infection 0 neoplastic process 0 consolidation by autologous bone marrow transplant 720 recurrent bacteremia and sepsis 720 multiorgan failure 720 death 1440
Mixtral-8x7B-Instruct v0.1 annotation
years old 0 male 0 recently diagnosed with lepromatous leprosy -60 on treatment (rifampicin/clofazimine/dapsone) -60 presented to the hospital 0 complaints of abdominal distension 0 complaints of constipation 0 complaints of vomiting 0 complaints of 10-kg weight loss 0 evidence of peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan of his abdomen showed mural thickening of the terminal ileum 0 computed tomography scan of his abdomen showed significantly enlarged mesenteric lymph nodes 0 computed tomography scan of his abdomen showed mesenteric fat stranding 0 computed tomography scan of his abdomen showed intra-abdominal free fluid 0 showing a dramatic response to treatment 0 planned for consolidation by autologous bone marrow transplant 0 re-admitted to the medical ICU for severe sepsis and multiorgan failure 0 passed away around 6 months after his initial diagnosis with NHL 0 despite maintaining a remission status 0
Llama-3.1-8B-Instruct annotation
years old 0 male 0 admitted to the hospital 0 lepromatous leprosy -720 treatment with rifampicin/clofazimine/dapsone -720 abdominal distension -720 constipation -720 vomiting -720 weight loss -720 peripheral lymphadenopathy -720 distended abdomen -720 positive shifting dullness -720 mural thickening of the terminal ileum -720 enlarged mesenteric lymph nodes -720 mesenteric fat stranding -720 intra-abdominal free fluid -720 abdominal granulomatous infection or neoplastic process -720 autologous bone marrow transplant -720 recurrent bacteremia -720 sepsis -720 multiorgan failure -720 passed away -720
Appendix D Additional Performance Analysis
Here we report performance plots for the sepsis-10 ablation analysis and the i2m4 dataset (main analysis and ablation), analogous to Figure 2. Figure D.1 displays the performance of the Llama 3.3 ablations on sepsis-10. In terms of match rate, the ablations perform within 5 percent of each other except for the 0-shot and Interval plus Typing (Int+Type) methods. In the temporal assessment figures we see that, for example, the 0-shot method shows promising temporal performance characteristics, but recall that it suffers from much lower event match rates. The Interval plus Typing method (Int+Type) shows a similar pattern. Requesting the interval alone (Interval) results in worse time discrepancy characteristics as compared to the main approach.
Figure D.1: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the sepsis-10 ablations. Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
Figures D.2 and D.3 displays the performance of the models and the Llama 3.3 ablations on the i2m4 dataset. The match rates and concordances are visibly lower in the i2m4 dataset than in sepsis-10. Comparing across methods, Llama 3.3 visibly outperforms the Llama 3.1 models in both measures. Despite the lower match rates and concordances, the time discrepancies in i2m4 were comparatively stronger (AULTC of 0.703). In terms of time discrepancy, Llama 3.1 70B demonstrate smaller time discrepancies, but at the cost of an event match rate of only 30 percent. In the ablation analysis of the i2m4 dataset, we observe that the Interval technique performs very similarly to the Llama 3.3 version (unlike in the sepsis-10 dataset, where it underperforms markedly with respect to time discrepancy).
Figure D.2: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the i2m4 dataset. Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
Figure D.3: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the i2m4 ablations. Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
Appendix E Error Analysis
E.1 Error Analysis of Event Match Rate
To investigate the discrepancies in event match rate between the i2m4 and sepsis-10 datasets, we analyzed the datasets’ structure and examined model performance using Llama 3.3 as a representative model, as the pattern of differences observed across datasets was consistent for all models.
Overview of Clinical Event Identification.
For the i2m4 dataset, manual annotations by clinicians identified an average of $195\pm 94$ clinical events per report across $20$ case reports, compared to $80\pm 28$ clinical events per report for sepsis-10 across $10$ case reports. Llama 3.3 identified an average of $120\pm 61$ clinical events for i2m4 and $78\pm 24$ clinical events for sepsis-10. These results highlight a systematic bias of under-identification of clinical events in i2m4 compared to sepsis-10 by Llama 3.3.
Additionally, the i2m4 dataset contains 2.3×2.3$\times$2.3 × more clinical events per report on average than sepsis-10, as annotated by the clinician, with a much larger variance in event counts across reports (Figure E.1). This heterogeneity reflects structural differences in the datasets: i2m4 reports often include a large number of structured entries, whereas sepsis-10 reports are only made-up of free text. Llama 3.3 over-identified clinical events (i.e., predicted more events than the manual annotators) in $15\%$ of i2m4 reports, while under-identifying events in $85\%$ of i2m4 reports. For sepsis-10, the model showed a more balanced distribution, over-identifying and under-identifying events in $50\%$ of reports each (Table E.1). These patterns suggest that the higher density and variability of clinical events in i2m4 pose a greater challenge for Llama 3.3.
Figure E.1: Comparison of Manual Annotation Clinical Event Counts and Llama 3.3 Model Predicted Event Counts: Scatter plot illustrating the variability between manual annotations and Llama 3.3 predictions across reports in two datasets, i2m4 (blue circles) and sepsis-10 (orange crosses). The red dashed line indicates the ideal one-to-one correspondence where the number of events per report matches exactly between manual annotations and model predictions.
Dataset % Reports Over-Identified % Reports Under-Identified
i2m4 15% 85%
sepsis-10 50% 50%
Table E.1: Proportion of reports where Llama 3.3 over- or under-identified clinical events compared to manual annotations. i2m4 shows a strong bias toward under-identification, whereas sepsis-10 has a more balanced distribution.
Dataset Structure Differences.
The differences in dataset structure between i2m4 and sepsis-10 are also evident in the length of clinical events. As shown in Table E.2, the manually annotated clinical events in i2m4 are significantly shorter, averaging $10\pm 7$ characters, compared to $28\pm 16$ characters in sepsis-10. This reflects the structured nature of i2m4 reports, which often include tables with abbreviations (e.g., “K” for Potassium, “Na” for Sodium). In contrast, sepsis-10 reports are written in free text and provide full terms and descriptions, avoiding the abbreviation issue entirely. The absence of abbreviations in sepsis-10 reduces variability in event representation and ensures greater alignment between manual annotations and Llama 3.3 predictions.
Dataset Manual Annotator Llama 3.3
i2m4 $10\pm 7$ $17\pm 14$
sepsis-10 $28\pm 16$ $21\pm 13$
Table E.2: Average length of clinical events in characters for both datasets. Manual annotations in i2m4 are significantly shorter due to frequent abbreviations and structured text from MIMIC table, while Llama 3.3 generates events of comparable length across datasets.
Llama 3.3-generated clinical events are of similar length across datasets, with averages of $17\pm 14$ characters for i2m4 and $21\pm 13$ characters for sepsis-10. This consistency suggests that Llama 3.3 applies similar generative behavior across datasets, often producing expanded forms of abbreviations in i2m4. While this generative behavior mitigates variability in Llama 3.3 outputs, it creates a mismatch with the shorter, more concise manual annotations in i2m4. This mismatch can then contribute to higher cosine distances between embeddings, lowering the event match rate for i2m4.
Event Match Rate Insights.
The event match rate, calculated as the proportion of manual clinical events correctly matched by Llama 3.3 predictions based on the sentence transformer embedding cosine distance threshold of 0.1, is significantly lower for i2m4 ($0.404$) than for sepsis-10 ($0.753$). It should be noted that over-identification of clinical events by Llama 3.3 does not impact this metric, as in calculating the event match rate, we only consider events identified by the manual annotators. The observed difference can be attributed to two main factors:
  • Under-Identification Bias in i2m4: When we adjust for under-identification by excluding manual clinical events for which Llama 3.3 did not identify a counterpart, the average match rate for i2m4 increases from $0.404$ to $0.752$, aligning more closely with sepsis-10’s match rate that increased from $0.753$ to $0.857$. Figure E.2 illustrates the match rate distributions across reports before and after the adjustment, highlighting the lower and more variable match rates in i2m4 compared to sepsis-10. This demonstrates that under-identification is the primary driver of the lower match rate in i2m4.
  • Dataset Structure Differences: The structured nature of i2m4 reports introduces variability in the manual annotations due to the frequent use of abbreviations and table-like formats. In contrast, the free-text nature of sepsis-10 ensures more consistent and complete representations of clinical events. While Llama 3.3 generates events of similar length across datasets, the expanded text spans of the abbreviations it generates in i2m4 exacerbate the mismatch with the manual annotations, leading to higher cosine distances and a lower event match rate.
Figure E.2: Distributions of Event Match Rates Across Reports Before and After Adjustment: (a) Unadjusted Match Rate: Histogram and density plots of event match rates between manual annotations and Llama 3.3 predictions for reports in i2m4 (blue) and sepsis-10 (orange). The lower and more variable match rates in i2m4 highlight the impact of under-identification bias. (b) Adjusted Match Rate: Histogram and density plots showing event match rates after excluding manual clinical events without Llama 3.3 counterparts. The adjustment improves the match rates in i2m4, demonstrating that under-identification is the primary cause of its lower initial match rates compared to sepsis-10.
These findings demonstrate that the lower event match rate for i2m4 is at least partially driven by a systematic under-identification of clinical events by Llama 3.3 in this dataset, compounded by structural differences of reports in i2m4 versus sepsis-10.
E.2 Error Analysis of Temporal Predictions
Across all models, i2m4 exhibits a lower concordance index (c-index) compared to sepsis-10. To better understand this discrepancy, we closely examine the performance of Llama 3.3, which shows similar trends in event ordering inconsistencies and a concordance-index drop from 0.93 to 0.75.
In i2m4, we observe higher variance in predicted event timestamps, with some extreme deviations from clinician annotations. To quantify the impact of these outliers, we apply an outlier filtering approach based on the interquartile range ($IQR$). The $IQR$ represents the middle 50% of time differences, defined as the range between the first quartile ($Q1$, $25^{th}$ percentile) and the third quartile ($Q3$, $75^{th}$ percentile). Predictions falling outside $Q1-1.5\times IQR$ or $Q3+1.5\times IQR$ are removed, ensuring that event ordering is evaluated on a more stable subset of predictions.
Figure E.3: Effect of Outlier Filtering on c-index in i2m4. Distribution of per-report c-index before and after removing extreme time mispredictions using the interquartile range (IQR). Filtering improves the average c-index by reducing the impact of large temporal deviations.
After filtering, the average c-index for i2m4 improves from 0.75 to 0.80, suggesting that extreme time mispredictions were an important factor in lowering event ranking consistency. The shift in per-report c-index distribution after filtering further supports this observation as shown in Figure E.3. However, despite this improvement, i2m4’s c-index remains lower than sepsis-10, indicating that additional challenges, such as increased annotation complexity or structural differences in event distributions mentioned in the previous section, contribute to its overall lower performance.
E.3 Clinical Event Type Categorization Process
Event Type Category Generation.
To systematically categorize clinical events identified in case reports from two datasets, sepsis-10 and i2m4, we employed a two-step methodology. The goal was to classify extracted clinical events into a predefined set of event type categories to use for further analysis of our results. Given i2m4’s data protection provisions only allowed for local analysis, we used sepsis-10 to determine the event type categories, as it was derived from PubMed Open Access case reports.
In the first step, event type categories were generated using a language model. Since the appropriate categories were not known a priori, we utilized OpenAI’s o3-mini, to analyze all manually annotated clinical events from the sepsis-10 dataset and derive six distinct event type categories. The model was prompted to examine the extracted events and construct a comprehensive yet non-overlapping set of categories that could effectively organize clinical events within the dataset. The full prompt used for this categorization process is as follows:
You are a physician and a machine learning scientist specializing in clinical event extraction. Your goal is to categorize extracted clinical events into exactly six distinct and mutually exclusive event categories to ensure clarity and consistency.
Constraints: Each clinical event must belong to exactly one category. The categories must be coarse enough to prevent ambiguity. The sixth category is reserved for “Other or Unknown.” The first five categories should comprehensively cover the key types of clinical events. The categories must be mutually exclusive, meaning an event should unambiguously fit into only one category.
Task: Identify six well-defined, non-overlapping event categories that best classify the extracted clinical events. Format the output as a comma-separated table with two columns: Event Type Category: The name of the category. Category Number: An integer from 0 to 5, with “Other or Unknown” always assigned to 5.
[Category Name], 0
[Category Name], 1
[Category Name], 2
[Category Name], 3
[Category Name], 4
Other or Unknown, 5
Now, based on the extracted clinical events below, generate six non-overlapping event categories and format the output as specified:
[…]
The resulting six event type categories, along with their assigned numeric labels, were:
  • Patient Background and Medical History (0) – Events describing patient demographics, prior medical conditions, risk factors, and relevant history before the clinical presentation.
  • Clinical Presentation and Examination Findings (1) – Events capturing the patient’s initial symptoms, subjective complaints, and observable clinical signs from physical examination.
  • Diagnostic Testing and Results (2) – Events related to diagnostic investigations, including laboratory tests, imaging studies, and other diagnostic assessments, along with their results.
  • Clinical Management and Interventions (3) – Events detailing treatments and medical interventions, including pharmacological therapies, surgical procedures, and other clinical management strategies.
  • Clinical Course, Outcomes, and Follow-up (4) – Events describing the progression of the condition, response to treatment, complications, recovery, prognosis, and post-treatment follow-up.
  • Other or Unknown (5) – Events that do not clearly fit into the defined categories or contain insufficient information to be classified.
Event Type Category Assignment.
Once these categories were established, they were used to systematically classify each identified clinical event from both sepsis-10 and i2m4. The classification was performed separately for clinical events identified through manual annotation and those identified using Llama 3.3. To maintain compliance with IRB regulations, all event classifications, including those for i2m4, were executed on a local machine to prevent data transfer to external servers.
For this classification step, we used Llama 3.3 to assign each extracted clinical event to one of the six predefined event type categories. The model was provided with a prompt that explicitly listed all event type categories and required it to assign each event to the most appropriate category. The classification was performed for both manually annotated events and events identified by Llama 3.3 across the sepsis-10 and i2m4 datasets. The exact prompt used for this classification task is as follows (the green text includes the few-shot examples provided to the model):
You are a medical professional. You are tasked with categorizing clinical events extracted from case reports.
Assign the following clinical event to one of these categories.
The categories are:
Patient Background and Medical History: 0,
Clinical Presentation and Examination Findings: 1,
Diagnostic Testing and Results: 2,
Clinical Management and Interventions: 3,
Clinical Course, Outcomes, and Follow‐up: 4,
Other or Unknown: 5
The categories are defined as follows:
Patient Background and Medical History: (Includes demographic details, prior medical diagnoses, past surgical histories, medication use as part of chronic history, and other baseline background information.),
Clinical Presentation and Examination Findings: (Includes symptoms at presentation, physical exam findings—including vital signs, neurological scores that reflect exam observations—and other immediate clinical observations.),
Diagnostic Testing and Results: (Includes all imaging studies, laboratory tests, diagnostic procedures and their reported findings, and formal diagnostic conclusions reached via workup.),
Clinical Management and Interventions: (Includes all treatments, procedures, medications administered acutely, operations, supportive care measures, and decisions/interventions intended to alter the patient’s condition.),
Clinical Course, Outcomes, and Follow‐up: (Includes statements about change in clinical status, response to treatment, complications, transitions in care, recovery, discharge, and long‐term outcomes.),
Other or Unknown: (For events that do not clearly fit into any of the above five categories.)
For example, here is a list of clinical event text and the corresponding category:
Examples:
“60-year-old female” → 0
“history of atrial fibrillation” → 0
“weighed 95 kg” → 0
“Impaired consciousness” → 1
“high-grade fever” → 1
“persistently high-temperature spikes” → 1
“Head CT” → 2
“no hepatitis A” → 2
“stage IV lymphoma” → 2
“successfully treated with fidaxomicin” → 3
“shifted to cefepime” → 3
“intravenous immunoglobulins for 5 days” → 3
“follow-up evaluations were recommended” → 4
“transferred to a geriatric medicine unit” → 4
“Discharge” → 4
“he” → 5
“confined, 5
“other symptoms” → 5
[…]
Respond with only the corresponding integer (0-5) from the list above. You have to pick only one category for each event. If there is no clear category, choose the category 5 that corresponds to “Other or Unknown” category. If there is more than one category, choose the category that you think is most relevant one. Do NOT include any extra text in your response. Do NOT show your thought process. Only provide the integer corresponding to the category and nothing else.
Format your response as:
Response: <integer >
[…]
Event Type Category Alignment and Match Rate Analysis.
To assess the consistency of event type categorization between the two annotators (manual annotator vs. Llama 3.3), we computed the event type category alignment rate for each dataset separately. This metric quantifies how often the manually annotated events and the Llama 3.3-extracted events—when determined to be aligned—are also assigned to the same event type category.
For each case report in the dataset, clinical events extracted by the manual annotator were aligned with those extracted by Llama 3.3 using an iterative matching approach based on the cosine similarity of their sentence-transformer embeddings. Aligned event pairs, which consist of one event from the manual annotator and one from Llama 3.3, were then categorized independently by Llama 3.3 into one of six predefined event type categories of the previous section.
The event type category alignment rate was computed by determining, for each aligned event pair, whether both events were assigned to the same category. If both events in an aligned pair received the same category label, the event type alignment was considered successful. For each dataset, the event type alignment rate was first calculated per case report by dividing the number of successfully aligned event type pairs by the total number of aligned pairs in that case report. These per-case alignment rates were then averaged across all case reports in the dataset to obtain the final alignment rate. Table E.3 reports the mean alignment rate along with the standard deviation for each dataset.
The event type category alignment rate was higher for the i2m4 dataset ($0.71\pm 0.05$) compared to the sepsis-10 dataset ($0.56\pm 0.05$). We hypothesize that this discrepancy is due to the absence of the “Other or Unknown” category in the sepsis-10 dataset, which likely introduced ambiguity in category assignments. Specifically, since no events in sepsis-10 were assigned to the “Other or Unknown” category, the model may have redistributed events that would have otherwise fallen into this category among the other five categories. This misalignment in categorization likely contributed to the lower agreement rate between the two annotators in the sepsis-10 dataset.
Dataset Event Type Alignment
i2m4 $0.71\pm 0.05$
sepsis-10 $0.56\pm 0.05$
Table E.3: Mean event type category alignment rate between the manual annotator and Llama 3.3 for each dataset. The alignment rate is computed as the proportion of aligned event pairs that were assigned to the same event type category. The reported values represent the mean alignment rate across case reports, with standard deviation indicated.
To further investigate the impact of event type categorization on event match rates, we analyzed the match rate separately for each of the six predefined event categories within both datasets. Figure E.4 displays the mean match rate for each event category, with error bars representing the standard deviation across case reports. The dashed horizontal lines indicate the overall mean match rate for all event types within each dataset (as seen in Table 2), providing a point of reference for comparison.
From the figure, it is evident that the event match rates exhibit variability across categories but do not show significant deviations that would suggest one category consistently performs better or worse than the others. For instance, in both datasets, categories such as “Diagnostic Testing and Results” and “Clinical Course, Outcomes, and Follow up” tend to achieve higher match rates compared to others, though the differences are not substantial enough to be statistically significant. Similarly, the match rates for “Clinical Management and Interventions” and “Patient Background and Medical History” fall closer to the overall dataset mean, suggesting these categories are neither particularly challenging nor exceptionally easy for alignment.
A notable observation is the “Other or Unknown” category, which is entirely absent in the sepsis-10 dataset. This absence introduces a unique challenge for categorization, as events that might have otherwise been assigned to this category are redistributed among the remaining categories. The i2m4 dataset, on the other hand, includes this category, and has a moderate event match rate for the event types in this category. This discrepancy could partially suggest that the style, structure, or format of the sepsis-10 dataset differs from that of i2m4. The absence of the “Other or Unknown” category in sepsis-10 could indicate a more standardized or narrowly focused structure, where fewer ambiguous or unclassifiable events are present, whereas i2m4 may include a broader range of events requiring such a fallback category. This difference in dataset characteristics could contribute to the observed variability in event match rates, as the annotation and alignment processes are inherently influenced by the dataset’s structure and complexity.
Figure E.4: Mean event match rate across predefined event type categories for the i2m4 and sepsis-10 datasets. Bars represent the mean match rate for each event type category, calculated as the proportion of aligned event pairs (between the manual annotator and Llama 3.3) with cosine similarity less than the matching threshold. Error bars indicate the standard deviation of match rates across case reports. The dashed lines represent the overall mean match rates for each dataset across all categories (From Table 2: sepsis-10: 0.753 and i2m4: 0.404) The “Other or Unknown” category is absent in sepsis-10, resulting in no match rate for this category in that dataset.
Overall, the analysis suggests that event match rates are relatively stable across event categories within each dataset. The differences in overall match rates between i2m4 and sepsis-10 are therefore more likely attributable to dataset-specific characteristics and under-identification bias that was explained in the previous section. Future work could explore more granular metrics, such as examining the semantic content of events within each category, to better understand the nuances of these event match rate discrepancies.

### v2

@@
 Our work characterizes the ability of LLMs to time-localize clinical findings in text, illustrating the limitations of LLM use for temporal reconstruction and providing several potential avenues of improvement via multimodal integration.
+keywords:
+Clinical Time Series, Sepsis
 1 Introduction
-The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria (Kyriazopoulou et al., 2021) and phenotyping studies (Seymour et al., 2019).
+The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria [1] and phenotyping studies [2] .
 In the critical care literature, sepsis is seen as a heterogeneous disease, a pathway arrived upon from a multitude of infectious origins, anatomical sites, and comorbid profiles.
-There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems (Henry et al., 2022), variation in prediction utility (Kamran et al., 2024), and endotype characterization (Noroozizadeh et al., 2023).
-Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis (Moldwin et al., 2021).
+There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems [3] , variation in prediction utility [4] , and endotype characterization [5] .
+Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis [6].
 A more complete alternative is the discharge summary, or more generally, the case report, in which a clinician documents the clinical findings pertinent to the case and care management.
@@
 At present, the (semi-)public critical care repositories at the research frontier contain high resolution structured signals and low resolution textual signals.
-While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts (Sun et al., 2013) rather than on event timing.
-Those that have focused on timing (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024) used excerpted summaries and relatively small sample sizes.
+While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts [7] rather than on event timing.
+Those that have focused on timing [8, 9] used excerpted summaries and relatively small sample sizes.
 In order to train sepsis models that can consider sources of information references in free-text and that have fine time-granularity, we seek to generate larger samples of textual time series.
@@
 We augment existing clinical annotations with our own on PMOA, and contrast LLM annotations with our own and that of Frattallone et al.
-(2024).
+(2024) [9] .
 To provide checks on their performance at scale, we consider a larger subsample for inter-LLM annotator agreement, and show the performance levels across several model sizes, which provides valuable context for cost-based analyses.
-Finally, these assessments provide evidence of the quality of the LLM annotations and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
+Finally, these assessments provide evidence of the LLM annotations quality and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
 1.1 Related work
-Numerous critical care and sepsis datasets are available (semi-)publicly, e.g., Thoral et al.
-(2021); Pollard et al.
-(2018), including several with de-identified clinical discharge summaries: MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023).
-Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data (Johnson et al., 2018), although structured data is often highly missing, leading to suboptimal phenotypes (Seinen et al., 2025).
-Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models (Kamran et al., 2024).
-Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024).
-In the case of Leeuwenberg and Moens (2020), the event spans are already given, and focus on an excerpt of the full discharge summary, so we focus our direct comparison on that of Frattallone-Llado et al.
-(2024).
-In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept (Uzuner et al., 2011), which allows for greater specificity of the event in the absence of additional context.
-Additionally, as compared to Frattallone-Llado et al.
-(2024), we focus on using text alone for clinical time series construction, which is necessary for the Pubmed Open-Access Subset since no complementary data source is available.
-Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization (Van Veen et al., 2024).
-In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models (Jeong et al., 2024).
-In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to the performance of the model.
+Numerous critical care and sepsis datasets are available (semi-)publicly [10, 11] , including several with de-identified clinical discharge summaries: MIMIC-III [12] and MIMIC-IV [13] .
+Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data [14] , although structured data is often highly missing, leading to suboptimal phenotypes [15].
+Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models [4] .
+Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts [8, 9] .
+In Leeuwenberg et al.
+[8] , the event spans are already given, and they focus on an excerpt of the discharge summary.
+Hence, we focus our direct comparison on that of Frattallone et al [9] .
+In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept [16] , which allows for greater specificity of the event in the absence of additional context.
+Additionally, as compared to Frattallone et al.
+[9] , we focus on using text alone for clinical time series construction, which is necessary for PMOA since no complementary data source is available.
+Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization [17] .
+In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models [18] .
+In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to their performance.
 1.2 Contributions
 Our work highlights the following contributions: (1) we present the first textual time series corpus for sepsis, comprising 2,139 open-access case reports identified as having sepsis; (2) we develop an analytic pipeline to identify Sepsis-3 from clinical text and to extract time-associated clinical findings; (3) we annotate time-associated clinical findings on a random subset of the case reports; (4) we develop and adapt evaluation criteria for textual time series, where matched clinical findings are not expected to be identical text strings; and, (5) in multiple datasets, we validate the performance of our pipeline, showing strong annotation performance compared to manual annotation and improved temporal results compared to existing methods.
-1.3 Generalizable Insights
-Access to the best temporal information available is critical in sepsis prediction tasks and has major implications in clinical decision making and early warning.
-The analysis pipeline we validate and corpus we introduce enable the use of textual data with temporal granularity, expanding the available features for use in a multitude of tasks.
-For example, the recent investigation in JAMA on vasopressin treatment policies using offline reinforcement learning uses hourly structured data in its learning, its reweighting analyses, and its validations (Kalimouttou et al., 2025).
-Access to textual features would augment the information that could affect its predictions, recommendations, and its adjustment for measured confounding, which may alter and refine its main findings.
-By adopting or extending our approach, the temporal resolution of the clinical event findings available only in text could be used and accounted for.
-As observed in Seinen et al.
-(2025), this may be a large fraction of concepts, with only 13% of textual events identified in structured form in their study.
 2 Methods
 We define the task specification, relevant terms, the extraction and annotation procedure, and the assessments.
-Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3.
-We define the term textual time series as a list of clinical findings each with an associated timestamp (which may be absolute or relative to time of case presentation) pertaining to an individual.
-A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
 i2m4
@@
 No 5 2
-Figure 1: Flowchart of the PMOA T2S2 corpus construction pipeline (left).
-Confusion matrices (right) for Sepsis-3 phenotype in i2m4 (top) and sepsis-100 (bottom) datasets, with preferred annotator in column position.
-Compared to clinical concept annotations (Uzuner et al., 2011; Sun et al., 2013), we choose to expand the text span determination to increase the specificity of the meaning attached to the finding.
-We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications (Uzuner et al., 2011):
+Figure 1: PMOA T2S2 pipeline (left) and Sepsis-3 confusion matrices (right)
+Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3.
+We define a textual time series as a list of clinical findings, each with a timestamp (absolute or relative to case presentation) for an individual.
+A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
+Compared to clinical concept annotations [16, 7] , we choose to expand the text span determination to increase the specificity of the meaning attached to the finding.
+We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications [16] :
   1. 1.
@@
 If the patient has a “history of smoking”, the start time is fully ambiguous, so the time of perception is the time that smoking is assessed, e.g., at the start of the encounter, and the event is “history of smoking”, rather than “smoking”.
-To restate, our clinical finding is defined as a contiguous text span from the text corresponding to a temporally localizable event pertaining to the patient, modified for conjunction splitting and temporal perception as above.
+Our clinical finding is defined as a contiguous text span representing a temporally localizable patient event, adjusted for conjunction splitting and temporal perception.
 Given this specification, our subtasks were: case report detection, Sepsis-3 detection, clinical finding extraction, and timestamp determination.
@@
 2.1 Data
-We used two data sources: (1) the absolute timeline annotations from Frattallone-Llado et al. (2024), and (2) the Pubmed Open Access subset (PMC Open Access Subset, 2024) (see Table 1).
+We used two data sources: (1) the absolute timeline annotations [9] , and (2) the Pubmed Open Access subset [19] (see Table 2.1.1).
 2.1.1 i2m4 Dataset
-The absolute timeline annotations (Frattallone-Llado et al., 2024) came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts.
+The absolute timeline annotations [9] came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts.
 For our analysis, we used the lower bound of the interval (specified as a mean time) as it denotes the time when the finding may first be observed.
 We used only the probabilistic annotations, since the bounded annotation lower bound could precede the actual occurrence of the finding.
-Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated in the same way)—hence we call this dataset “i2m4”.
-Table 1: Dataset summary. Bold cells indicate contributed labeling
-Corpus i2m4 Sepsis-10   Sepsis-100
-N (unlabeled) 310; 331,794 2,139 2,139
-N (labeled) 20 10 100
-Source i2b2; MIMIC IV PMOA PMOA
-Note type discharge summary case report case report
-Labeler clinicians clinicians O1-preview
-Phenotyper clinicians clinicians QwQ-32B-Preview
-Methods O1-preview O1-preview
-DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1
-L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct
-L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct
-Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B
- Instruct v0.1  Instruct v0.1  Instruct v0.1
-L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct
-Ablations No role No role No role
-Base: No conjunction No conjunction No conjunction
-L3.3 70B Instruct  instruction  instruction  instruction
-No few-shot prompting No few-shot prompting No few-shot prompting
-Sensitivity analyses With upper bound With upper bound With upper bound
-With upper bound + With upper bound + With upper bound +
- i2b2 event typing  i2b2 event typing  i2b2 event typing
+Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated similarly)—hence we call this dataset “i2m4”.
+[…]
+Dataset summary.
+Bold cells indicate contributed labeling Corpus i2m4 Sepsis-10   Sepsis-100 N (unlabeled) 310; 331,794 2,139 2,139 N (labeled) 20 10 100 Source i2b2; MIMIC IV PMOA PMOA Note type discharge summary case report case report Labeler clinicians clinicians O1-preview Phenotyper clinicians clinicians QwQ-32B-Preview Methods O1-preview O1-preview DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1 L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B  Instruct v0.1  Instruct v0.1  Instruct v0.1 L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct Ablations No role No role No role Base: No conjunction No conjunction No conjunction L3.3 70B Instruct  instruction  instruction  instruction No few-shot prompting No few-shot prompting No few-shot prompting Sensitivity analyses With upper bound With upper bound With upper bound With upper bound + With upper bound + With upper bound +  i2b2 event typing  i2b2 event typing  i2b2 event typing
 For Sepsis-3 determination, we had a clinician review the discharge summaries and assign each case as meeting the Sepsis-3 definition.
@@
 2.1.2 PMOA (sepsis-10, sepsis-100) Dataset
-The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications (PMC Open Access Subset, 2024).
+The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications [19] .
 Because the majority of these publications are not case reports, we built an extraction pipeline to recover critical care case reports, i.e., a clinical research informatics critical care cohort.
@@
 ?old.
-We contrasted the output of the query with PubMed metadata for case reports and found our regular expression approach had substantially higher specificity during inspection of a sample of 20 random reports.
-To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||||septic) and (critical||||intensive) care.
+We compared the query output to PubMed case report metadata and found our regex approach had much higher specificity in a sample of 20 random reports.
+To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||septic) and (critical||intensive) care.
 Then, we used the an LLM query to extract sepsis cases, as well as the number of cases present, and the age and gender.
@@
 Then the output should look like
-18 years old |||| 0
-male |||| 0
-admitted to the hospital |||| 0
-fever |||| -72
-rash |||| -72
-acne |||| -672
-minocycline |||| -672
-increased WBC count |||| 0
-eosinophilia|||| 0
-systemic involvement|||| 0
-diffuse erythematous or maculopapular eruption|||| 0
-pruritis |||| 0
-DRESS syndrome |||| 0
-fever persisted |||| 0
-rash persisted |||| 0
-discharged |||| 24
+18 years old $|$ 0
+male $|$ 0
+admitted to the hospital $|$ 0
+fever $|$ -72
+rash $|$ -72
+acne $|$ -672
+minocycline $|$ -672
+increased WBC count $|$ 0
+eosinophilia|| 0
+systemic involvement|| 0
+diffuse erythematous or maculopapular eruption|| 0
+pruritis $|$ 0
+DRESS syndrome $|$ 0
+fever persisted $|$ 0
+rash persisted $|$ 0
+discharged $|$ 24
 […]
@@
 For the i2m4 dataset, we also compared against a BERT-based baseline (BERT 5-CV), our adaption of the Frattallone et al.
-model (2024).
+model (2024) [9] .
 In their work, they transformed the temporal prediction task into multiclass classification and used a multilayer perceptron (MLP) head with a softmax and cross-entropy loss for training.
@@
 The regression task (predicting the lower bound) was optimized using L1 loss.
-We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span (Zhong and Chen, 2021) and removed the dropout layers inside the MLP heads, leaving only one at the final layer.
+We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span [20] and removed the dropout layers inside the MLP heads, leaving only one at the final layer.
 Regarding the timestamps used for prediction, while many events occurred near the admission time, a subset extended months or years away, complicating the design of a regressor capable of both spanning wide magnitudes and maintaining precision for the majority of the means close to the admission time.
@@
 We assessed the textual time series with the following methods: event match rate, concordance (c-index), median absolute time discrepancy, and log-time discrepancy.
-To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix B for pseudocode).
+To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix 2 for pseudocode).
 Specifically, the recursive best match was used to reduce erroneous surjective matches, i.e., multiple events from the predicted event list mapping to the same event in the reference list.
@@
 Then all matched events from the recursive calls were returned.
-To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) (Deka et al., 2022).
+To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) [21].
 Upon manual review of event matches, we selected option 3 with a cosine similarity cutoff of 0.1.
@@
 For each matched event $i$, compute the discrepancy $s_{i}=|t^{p}_{i}-t^{r}_{i}|$.
-then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours):
-$s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$
-Apply the log transformation:
-$x_{i}=\log(1+s^{\prime}_{i})$
-Sort $x_{i}$ in ascending order:
-$x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$
-The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$:
-$F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$
-where $\mathbf{1}$ is the indicator function.
+then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours): $s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$.
+Apply the log transformation: $x_{i}=\log(1+s^{\prime}_{i})$ Sort $x_{i}$ in ascending order: $x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$.
+The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$: $F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$ where $\mathbf{1}$ is the indicator function.
 Then, we get the following boundary values: $F(0)\geq 0$ (depending on the smallest discrepancy) and $F(\log(1+S_{\text{max}}))=1$.
@@
 With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area.
-We provide additional remarks regarding the need for explicit bound specification of the log-time CDF and AULTC and the non-convexity of the average log discrepancy in Appendix A.
+Additional remarks on the necessity of explicit bounds for the log-time CDF and AULTC, as well as the non-convexity of the average log discrepancy, are provided in Appendix References.
 2.4 Annotation Quality Review
@@
 The annotator also placed the annotations into quality levels of Excellent, Good, Acceptable, and Poor.
-The guideline for placing the annotations into each quality level was:
+The guideline for placing the annotations into each quality level:
 Excellent: High quality extraction. May contain up to 1 clear error in extraction or timing; extracts all remaining important findings; ordering and timing are plausible.
@@
 The annotator was asked to document pertinent findings and errors in free text.
-Table 2: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets.
-i2m4 sepsis-10 sepsis-100
-Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
-O1-preview 0.76 0.932 22 0.796
-DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801
-L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
-L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790
-Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737
-L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705
-BERT 5-CV 0.704 34 0.520
+[…]
+Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets.
+i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC O1-preview 0.76 0.932 22 0.796 DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801 L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790 Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737 L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705 BERT 5-CV 0.704 34 0.520
 3 Results
@@
 The sepsis-10 manual annotations are of 10 files: PMC3075162, PMC4778217, PMC6238558, PMC7576661, PMC9136437, PMC9355071, PMC9552858, PMC10370906, PMC10556716, and PMC10629858, resulting in 801 reference events.
-An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix C.
+An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix .1.
 3.1 Phenotyping Results
@@
 3.2 Annotation Results
-The results in Table 2 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets.
-In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list (Frattallone-Llado et al., 2024)), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding.
+The results in Table 2.4 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets.
+In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list [9]), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding.
 Among the matches, Llama 3.3 and DeepSeek R1 performed best for time predictions (trading off event matching with temporal performance), and substantially outperforming the baseline BERT model, particularly with respect to time discrepancy.
@@
 In contrast to concordances of 0.707 in i2m4, in sepsis-10, O1-preview and Llama 3.3 were excellent time-ordering annotators, both with concordances of 0.932.
-Table 3: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for i2m4, sepsis-10, and sepsis-100.
-i2m4 sepsis-10 sepsis-100
-Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
-L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
-Ablations
-No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789
-No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793
-Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795
-Augments
-Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791
-Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
+[…]
+Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for our 3 datasets.
+i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 Ablations No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789 No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793 Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795 Augments Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791 Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
 To examine the results in greater detail, we show figures of the cumulative distribution function of event matches and concordance boxplots (Figure 2) for the sepsis-10 dataset.
-When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event that the prediction method did not.
-The event match rates in Table 2 correspond to the match rate at the cosine distance threshold in Figure 2.
-One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positive matches, which degrade the temporal performance characteristics.
-The concordance box plots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower, which is a consistent finding throughout our analysis.
+When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event the prediction method did not.
+The event match rates in Table 2.4 correspond to the match rate at the cosine distance threshold in Figure 2.
+One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positives, which degrade temporal performance characteristics.
+The concordance boxplots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower—a consistent finding throughout our analysis.
 To better visualize the time discrepancies, we plot cumulative time discrepancy plots (Figure 2) that show, for example, that 44% of the O1-preview annotations are exact time matches and that 80% are within 1 week of the label timestamp.
@@
 Problematically, for labeled events more than 1 year from presentation, only O1-preview and DeepSeek R1 ever identify the label time to within 1 year exclusive (Figure 2, right).
-The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3).
+The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3.2).
 Our original prompt appears to be the strong performer overall, showcasing the value in particular of few-shot prompting, but also the value of the technical specification of how to treat conjunctions and the modest benefit of role-playing.
@@
 The performance of the typing method (Interval + Type) is further degraded.
-Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix D.
-To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix E.
+Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis.
+To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix .5.
 This analysis provides additional insight into the factors contributing to variations in model performance for event match rate and c-index across datasets.
-Figure 2: Event match cumulative distribution function (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
+Figure 2: Event match CDF (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
 3.3 Annotation Quality Results
@@
 While the Llama 3.3 annotations are found to be acceptable a high percentage of the time (90%), the other Llama and Mistral approaches fail to achieve Acceptable performance in excess of 50%.
-Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2 and Figure 2 (left and middle right)).
+Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2.4 and Figure 2 (left and middle right)).
 Our clinical reviewer commented that the Llama 3.3 discrepancies for large time horizons, i.e., years, were large, which is corroborated in Figure 2 (right, bottom subpanels).
@@
 However, the event can also be considered as “history of 10-kg weight loss” via the “history of” guideline with time 0 (what L3.3 70B annotates).
-Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3 and Figure D.1).
+Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3.2 and Figure .3).
 Events requiring temporal reasoning or composition.
@@
 In PMOA samples sepsis-10 and sepsis-100, performance levels were strong with respect to clinician annotations (sepsis-10) and inter-LLM agreement (sepsis-100).
-Table 4: Sepsis-10 annotation quality review
-Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable
-Manual 1.2 0.9 1 1 1
-O1-preview 1.5 0.6 0.9 0.9 1 1
-DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8
-L3.3 70B 3.8 0 0 0 0.2 0.9
-L3.1 70B 4.8 0 0 0 0.2 0.5
-Mixtral 8x7B 5.4 0 0 0 0.1 0.2
-L3.1 8B 6.5 0 0 0 0 0
+[…]
+Sepsis-10 annotation quality review Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable Manual 1.2 0.9 1 1 1 O1-preview 1.5 0.6 0.9 0.9 1 1 DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8 L3.3 70B 3.8 0 0 0 0.2 0.9 L3.1 70B 4.8 0 0 0 0.2 0.5 Mixtral 8x7B 5.4 0 0 0 0.1 0.2 L3.1 8B 6.5 0 0 0 0 0
 There are several limitations to our work.
@@
 Also, the amount and type of information in such case reports may vary.
-Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when you mix intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales will not dominate the performance measure (because of the time discrepancy log scaling).
+Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when mixing intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales do not dominate the performance measure (because of the time discrepancy log scaling).
 Another generalizability concern is that for some conditions, less temporal information will be present.
@@
 Future work could further characterize the findings, with additional or improved augmentations to the current system.
-Although our augmentation strategies did not improve overall system performance (Table 3), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation.
+Although our augmentation strategies did not improve overall system performance (Table 3.2), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation.
 Our hope is that this corpus, or a future version of it, will be used for downstream applications in sepsis, such as causal modeling, forecasting, and multimodal alignment.
@@
 This work utilized the computational resources of the NIH HPC Biowulf cluster.
-Appendix A Log-Time Cumulative Distribution Function
-Recall the log-time cumulative distribution function is given as follows:
-$F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$
-where $\mathbf{1}$ is the indicator function.
-We compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
-AULTC $=\frac{1}{\log(1+S_{\text{max}})}\Bigg{[}\sum_{i=1}^{k}(x_{(i)}-x_{(i-1)})\frac{i}{k}+(\log(1+S_{\text{max}})-x_{(k)})\times 1\Bigg{]}$
-where $x_{(0)}=0$.
-With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area.
-The code for the AULTC calculation is given in the function ecdf_auc2(…) in the Supplementary Materials.
-Remark A.1.
-The time unit and cutoff $S_{\text{max}}$ affect the AULTC calculation, so they must be specified when reporting the AULTC.
-Because of the $\log(1+\cdot)$ transformation (rather than the $\log(\cdot)$ transformation which is undefined for zero time-error discrepancy), the discrepancies are non-linearly shifted in the log scale.
-The non-linear shift adjusts the relative widths of the step function (particularly for small discrepancies) changing the area calculation.
-The cutoff $S_{\text{max}}$ affects the normalization factor.
-Therefore, these values should be chosen based on practicalities for the application.
-In our case, we chose hours at the time unit because the only sub-hour descriptions reported were several Apgar scores at the minute-level.
-Remark A.2.
-The average log-time discrepancy is non-convex.
-This can be seen from the observation that the $\log(1+\cdot)$ function is “spikier” than the L1 function.
-More formally, the log-time discrepancy has a higher curvature around zero error compared to the L1 loss.
-Consider the second derivatives of the loss functions with respect to the discrepancies $s=|t^{p}-t^{r}|$.
-For L1 loss $L_{1}(s)=s$, the second derivative $\frac{d^{2}L_{1}}{ds^{2}}=0$ for $s>0$.
-For the log-time loss, $L_{log}(s)=\log(1+s)$, the second derivative is $\frac{d^{2}L_{log}}{ds^{2}}=-\frac{1}{(1+s)^{2}}$.
-At $s=0$, the second derivative is $-1$, while the second derivative of the L1 loss is 00.
-To provide a simple counterexample, define the average log-time discrepancy is given by:
-$L(t^{p})=\frac{1}{k}\sum_{i=1}^{k}\log(1+\min(|t^{p}_{i}-t^{r}_{i}|,S_{\text{max}}))$
-To prove that this function is non-convex, we need to show that there exist $t^{p1},t^{p2}$ and $\lambda\in(0,1)$ such that
-$L(\lambda t^{p1}+(1-\lambda)t^{p2})>\lambda L(t^{p1})+(1-\lambda)L(t^{p2})$
-Consider the case where $k=1$, $t^{r}=0$, and $S_{\text{max}}=2$.
-The loss function is $L(t^{p})=\log(1+\min(|t^{p}|,2))$.
-Let $t^{p1}=0$ and $t^{p2}=3$.
-Choose $\lambda=0.5$.
-Then $\lambda t^{p1}+(1-\lambda)t^{p2}=0.5\times 0+0.5\times 3=1.5$.
-Now we evaluate the loss function at these points:
-$L(t^{p1})$ $=L(0)=\log(1+\min(|0|,2))=0$
-$L(t^{p2})$ $=L(3)=\log(1+\min(|3|,2))=$
-$\qquad\log(3)$
-$L(\lambda t^{p1}+(1-\lambda)t^{p2})$ $=L(1.5)=\log(2.5)$
-Now we check the convexity condition:
-$L(1.5)\leq 0.5L(0)+0.5L(3)$
-$\log(2.5)\leq 0.5\times 0+0.5\times\log(3)$
-$\log(2.5)\leq\log(\sqrt{3})$
-Since $2.5>\sqrt{3}\approx 1.732$, we have $\log(2.5)>\log(\sqrt{3})$. Therefore,
-$L(\lambda t^{p1}+(1-\lambda)t^{p2})>\lambda L(t^{p1})+(1-\lambda)L(t^{p2})$
-This proves that the average log-time discrepancy is non-convex. This remark illustrates the loss form, in case it is considered for optimization/model training rather than for assessment purposes.
-Appendix B Recursive Best Match Procedure
-We provide pseudocode for the best match procedure between two lists of strings (Algorithm B.1).
-For each list, we use the text order, i.e., the order of the events in the annotation files, to break embedding distance ties.
-We use the cosine similarity for the distance calculation using sentence transformer embeddings from S-PubMedBert-MS-MARCO.
-Algorithm B.1 Recursive Best Match
-Input :
-Two lists: ref (reference events) and pred (predicted events)
-Output :
-List of best-matching event pairs
-MatchEvents(ref, pred)
-if ref is empty or pred is empty then
-       return []
-end if
-[…]
-       foreach $p$ in pred do
-[…]
-[…]
-             end if
-[…]
-[…]
-[…]
-                   end if
-[…]
-[…]
-                   end if
-             end if
-       end foreach
-end foreach
-[…]
-Appendix C Example Annotation
-In this section we present an excerpt from PMC10629858, the most recently published case in the sepsis-10 dataset (Abu-Tineh et al., 2023).
-To illustrate the extraction task and performance characteristics of the LLM annotators, we contrast the manual annotations with that of O1-preview, Llama 3.3, and Llama 3.1 8B.
-Excerpt from Abu-Tineh et al. (2023):
-A 57-year-old man recently diagnosed with lepromatous leprosy was confirmed with skin biopsy and had been on treatment (rifampicin/clofazimine/dapsone) for 2 months before admission; he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss.
-On examination, the patient was vitally stable.
-He had evidence of peripheral lymphadenopathy with a distended abdomen and a positive shifting dullness.
-A computed tomography scan of his abdomen showed mural thickening of the terminal ileum with significantly enlarged mesenteric lymph nodes, mesenteric fat stranding, and intra-abdominal free fluid, suggesting abdominal granulomatous infection or neoplastic process.
-…
-The patient was planned for consolidation by autologous bone marrow transplant.
-Unfortunately, with the recurrent bacteremia and sepsis that accompanied the patient’s course due to his immunocompromised state, he was re-admitted to the medical ICU for severe sepsis and multiorgan failure and passed away around 6 months after his initial diagnosis with NHL, despite maintaining a remission status.
-Manual annotation
--year-old 0 man 0 diagnosed with lepromatous leprosy -1461 on treatment -1461 rifampicin -1461 clofazimine -1461 dapsone -1461 presented to the hospital 0 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 vitally stable 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan of his abdomen 0 mural thickening of the terminal ileum 0 significantly enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 re-admitted to the medical ICU 4383 severe sepsis 4383 multiorgan failure 4383 passed away 4383
-O1-preview annotation
-Recently diagnosed with lepromatous leprosy -1440 Skin biopsy confirmed lepromatous leprosy -1440 Started on rifampicin -1440 Started on clofazimine -1440 Started on dapsone -1440 -kg weight loss -168 -year-old man 0 Presented to the hospital 0 Complaints of abdominal distension 0 Complaints of constipation 0 Complaints of vomiting 0 Patient was vitally stable 0 Evidence of peripheral lymphadenopathy 0 Distended abdomen 0 Positive shifting dullness 0 Computed tomography scan of abdomen 0 Mural thickening of terminal ileum 0 Significantly enlarged mesenteric lymph nodes 0 Mesenteric fat stranding 0 Intra-abdominal free fluid 0 Planned for consolidation by autologous bone marrow transplant 2136 Responded well to chemotherapy 2280 Attained complete metabolic remission by positron emission tomography/computed tomography 2280 Re-admitted to medical ICU for severe sepsis and multiorgan failure 4360 Passed away around 6 months after his initial diagnosis with NHL 4368
-DeepSeek-R1-IQ1 annotation
-years old 0 male 0 lepromatous leprosy -1464 skin biopsy -1464 rifampicin -1464 clofazimine -1464 dapsone -1464 admitted to the hospital 0 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 vital stability 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan 0 mural thickening of the terminal ileum 0 enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 abdominal granulomatous infection or neoplastic process 0 plan for autologous bone marrow transplant 0 recurrent bacteremia 0 sepsis 0 readmission to medical ICU 0 severe sepsis 24 multiorgan failure 24 passed away 4032
-Llama-3.3-70B-Instruct annotation
-years old 0 male 0 lepromatous leprosy -720 treatment with rifampicin/clofazimine/dapsone -720 abdominal distension 0 constipation 0 vomiting 0 -kg weight loss 0 peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 mural thickening of the terminal ileum 0 enlarged mesenteric lymph nodes 0 mesenteric fat stranding 0 intra-abdominal free fluid 0 abdominal granulomatous infection 0 neoplastic process 0 consolidation by autologous bone marrow transplant 720 recurrent bacteremia and sepsis 720 multiorgan failure 720 death 1440
-Mixtral-8x7B-Instruct v0.1 annotation
-years old 0 male 0 recently diagnosed with lepromatous leprosy -60 on treatment (rifampicin/clofazimine/dapsone) -60 presented to the hospital 0 complaints of abdominal distension 0 complaints of constipation 0 complaints of vomiting 0 complaints of 10-kg weight loss 0 evidence of peripheral lymphadenopathy 0 distended abdomen 0 positive shifting dullness 0 computed tomography scan of his abdomen showed mural thickening of the terminal ileum 0 computed tomography scan of his abdomen showed significantly enlarged mesenteric lymph nodes 0 computed tomography scan of his abdomen showed mesenteric fat stranding 0 computed tomography scan of his abdomen showed intra-abdominal free fluid 0 showing a dramatic response to treatment 0 planned for consolidation by autologous bone marrow transplant 0 re-admitted to the medical ICU for severe sepsis and multiorgan failure 0 passed away around 6 months after his initial diagnosis with NHL 0 despite maintaining a remission status 0
-Llama-3.1-8B-Instruct annotation
-years old 0 male 0 admitted to the hospital 0 lepromatous leprosy -720 treatment with rifampicin/clofazimine/dapsone -720 abdominal distension -720 constipation -720 vomiting -720 weight loss -720 peripheral lymphadenopathy -720 distended abdomen -720 positive shifting dullness -720 mural thickening of the terminal ileum -720 enlarged mesenteric lymph nodes -720 mesenteric fat stranding -720 intra-abdominal free fluid -720 abdominal granulomatous infection or neoplastic process -720 autologous bone marrow transplant -720 recurrent bacteremia -720 sepsis -720 multiorgan failure -720 passed away -720
-Appendix D Additional Performance Analysis
-Here we report performance plots for the sepsis-10 ablation analysis and the i2m4 dataset (main analysis and ablation), analogous to Figure 2.
-Figure D.1 displays the performance of the Llama 3.3 ablations on sepsis-10.
-In terms of match rate, the ablations perform within 5 percent of each other except for the 0-shot and Interval plus Typing (Int+Type) methods.
-In the temporal assessment figures we see that, for example, the 0-shot method shows promising temporal performance characteristics, but recall that it suffers from much lower event match rates.
-The Interval plus Typing method (Int+Type) shows a similar pattern.
-Requesting the interval alone (Interval) results in worse time discrepancy characteristics as compared to the main approach.
-Figure D.1: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the sepsis-10 ablations.
-Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
-Figures D.2 and D.3 displays the performance of the models and the Llama 3.3 ablations on the i2m4 dataset.
-The match rates and concordances are visibly lower in the i2m4 dataset than in sepsis-10.
-Comparing across methods, Llama 3.3 visibly outperforms the Llama 3.1 models in both measures.
-Despite the lower match rates and concordances, the time discrepancies in i2m4 were comparatively stronger (AULTC of 0.703).
-In terms of time discrepancy, Llama 3.1 70B demonstrate smaller time discrepancies, but at the cost of an event match rate of only 30 percent.
-In the ablation analysis of the i2m4 dataset, we observe that the Interval technique performs very similarly to the Llama 3.3 version (unlike in the sepsis-10 dataset, where it underperforms markedly with respect to time discrepancy).
-Figure D.2: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the i2m4 dataset.
-Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
-Figure D.3: Event match cumulative distribution function (top left) and concordance box-plots (top right) for the i2m4 ablations.
-Time discrepancy from the manual annotation timestamps among matched events, overall (bottom left) and disaggregated by clinician annotator timestamp (time from presentation, bottom right).
-Appendix E Error Analysis
-E.1 Error Analysis of Event Match Rate
-To investigate the discrepancies in event match rate between the i2m4 and sepsis-10 datasets, we analyzed the datasets’ structure and examined model performance using Llama 3.3 as a representative model, as the pattern of differences observed across datasets was consistent for all models.
-Overview of Clinical Event Identification.
-For the i2m4 dataset, manual annotations by clinicians identified an average of $195\pm 94$ clinical events per report across $20$ case reports, compared to $80\pm 28$ clinical events per report for sepsis-10 across $10$ case reports.
-Llama 3.3 identified an average of $120\pm 61$ clinical events for i2m4 and $78\pm 24$ clinical events for sepsis-10.
-These results highlight a systematic bias of under-identification of clinical events in i2m4 compared to sepsis-10 by Llama 3.3.
-Additionally, the i2m4 dataset contains 2.3×2.3$\times$2.3 × more clinical events per report on average than sepsis-10, as annotated by the clinician, with a much larger variance in event counts across reports (Figure E.1).
-This heterogeneity reflects structural differences in the datasets: i2m4 reports often include a large number of structured entries, whereas sepsis-10 reports are only made-up of free text.
-Llama 3.3 over-identified clinical events (i.e., predicted more events than the manual annotators) in $15\%$ of i2m4 reports, while under-identifying events in $85\%$ of i2m4 reports.
-For sepsis-10, the model showed a more balanced distribution, over-identifying and under-identifying events in $50\%$ of reports each (Table E.1).
-These patterns suggest that the higher density and variability of clinical events in i2m4 pose a greater challenge for Llama 3.3.
-Figure E.1: Comparison of Manual Annotation Clinical Event Counts and Llama 3.3 Model Predicted Event Counts: Scatter plot illustrating the variability between manual annotations and Llama 3.3 predictions across reports in two datasets, i2m4 (blue circles) and sepsis-10 (orange crosses).
-The red dashed line indicates the ideal one-to-one correspondence where the number of events per report matches exactly between manual annotations and model predictions.
-Dataset % Reports Over-Identified % Reports Under-Identified
-i2m4 15% 85%
-sepsis-10 50% 50%
-Table E.1: Proportion of reports where Llama 3.3 over- or under-identified clinical events compared to manual annotations.
-i2m4 shows a strong bias toward under-identification, whereas sepsis-10 has a more balanced distribution.
-Dataset Structure Differences.
-The differences in dataset structure between i2m4 and sepsis-10 are also evident in the length of clinical events.
-As shown in Table E.2, the manually annotated clinical events in i2m4 are significantly shorter, averaging $10\pm 7$ characters, compared to $28\pm 16$ characters in sepsis-10.
-This reflects the structured nature of i2m4 reports, which often include tables with abbreviations (e.g., “K” for Potassium, “Na” for Sodium).
-In contrast, sepsis-10 reports are written in free text and provide full terms and descriptions, avoiding the abbreviation issue entirely.
-The absence of abbreviations in sepsis-10 reduces variability in event representation and ensures greater alignment between manual annotations and Llama 3.3 predictions.
-Dataset Manual Annotator Llama 3.3
-i2m4 $10\pm 7$ $17\pm 14$
-sepsis-10 $28\pm 16$ $21\pm 13$
-Table E.2: Average length of clinical events in characters for both datasets.
-Manual annotations in i2m4 are significantly shorter due to frequent abbreviations and structured text from MIMIC table, while Llama 3.3 generates events of comparable length across datasets.
-Llama 3.3-generated clinical events are of similar length across datasets, with averages of $17\pm 14$ characters for i2m4 and $21\pm 13$ characters for sepsis-10.
-This consistency suggests that Llama 3.3 applies similar generative behavior across datasets, often producing expanded forms of abbreviations in i2m4.
-While this generative behavior mitigates variability in Llama 3.3 outputs, it creates a mismatch with the shorter, more concise manual annotations in i2m4.
-This mismatch can then contribute to higher cosine distances between embeddings, lowering the event match rate for i2m4.
-Event Match Rate Insights.
-The event match rate, calculated as the proportion of manual clinical events correctly matched by Llama 3.3 predictions based on the sentence transformer embedding cosine distance threshold of 0.1, is significantly lower for i2m4 ($0.404$) than for sepsis-10 ($0.753$).
-It should be noted that over-identification of clinical events by Llama 3.3 does not impact this metric, as in calculating the event match rate, we only consider events identified by the manual annotators.
-The observed difference can be attributed to two main factors:
-  • Under-Identification Bias in i2m4: When we adjust for under-identification by excluding manual clinical events for which Llama 3.3 did not identify a counterpart, the average match rate for i2m4 increases from $0.404$ to $0.752$, aligning more closely with sepsis-10’s match rate that increased from $0.753$ to $0.857$.
-Figure E.2 illustrates the match rate distributions across reports before and after the adjustment, highlighting the lower and more variable match rates in i2m4 compared to sepsis-10.
-This demonstrates that under-identification is the primary driver of the lower match rate in i2m4.
-  • Dataset Structure Differences: The structured nature of i2m4 reports introduces variability in the manual annotations due to the frequent use of abbreviations and table-like formats.
-In contrast, the free-text nature of sepsis-10 ensures more consistent and complete representations of clinical events.
-While Llama 3.3 generates events of similar length across datasets, the expanded text spans of the abbreviations it generates in i2m4 exacerbate the mismatch with the manual annotations, leading to higher cosine distances and a lower event match rate.
-Figure E.2: Distributions of Event Match Rates Across Reports Before and After Adjustment: (a) Unadjusted Match Rate: Histogram and density plots of event match rates between manual annotations and Llama 3.3 predictions for reports in i2m4 (blue) and sepsis-10 (orange).
-The lower and more variable match rates in i2m4 highlight the impact of under-identification bias.
-(b) Adjusted Match Rate: Histogram and density plots showing event match rates after excluding manual clinical events without Llama 3.3 counterparts.
-The adjustment improves the match rates in i2m4, demonstrating that under-identification is the primary cause of its lower initial match rates compared to sepsis-10.
-These findings demonstrate that the lower event match rate for i2m4 is at least partially driven by a systematic under-identification of clinical events by Llama 3.3 in this dataset, compounded by structural differences of reports in i2m4 versus sepsis-10.
-E.2 Error Analysis of Temporal Predictions
-Across all models, i2m4 exhibits a lower concordance index (c-index) compared to sepsis-10.
-To better understand this discrepancy, we closely examine the performance of Llama 3.3, which shows similar trends in event ordering inconsistencies and a concordance-index drop from 0.93 to 0.75.
-In i2m4, we observe higher variance in predicted event timestamps, with some extreme deviations from clinician annotations.
-To quantify the impact of these outliers, we apply an outlier filtering approach based on the interquartile range ($IQR$).
-The $IQR$ represents the middle 50% of time differences, defined as the range between the first quartile ($Q1$, $25^{th}$ percentile) and the third quartile ($Q3$, $75^{th}$ percentile).
-Predictions falling outside $Q1-1.5\times IQR$ or $Q3+1.5\times IQR$ are removed, ensuring that event ordering is evaluated on a more stable subset of predictions.
-Figure E.3: Effect of Outlier Filtering on c-index in i2m4.
-Distribution of per-report c-index before and after removing extreme time mispredictions using the interquartile range (IQR).
-Filtering improves the average c-index by reducing the impact of large temporal deviations.
-After filtering, the average c-index for i2m4 improves from 0.75 to 0.80, suggesting that extreme time mispredictions were an important factor in lowering event ranking consistency.
-The shift in per-report c-index distribution after filtering further supports this observation as shown in Figure E.3.
-However, despite this improvement, i2m4’s c-index remains lower than sepsis-10, indicating that additional challenges, such as increased annotation complexity or structural differences in event distributions mentioned in the previous section, contribute to its overall lower performance.
-E.3 Clinical Event Type Categorization Process
-Event Type Category Generation.
-To systematically categorize clinical events identified in case reports from two datasets, sepsis-10 and i2m4, we employed a two-step methodology.
-The goal was to classify extracted clinical events into a predefined set of event type categories to use for further analysis of our results.
-Given i2m4’s data protection provisions only allowed for local analysis, we used sepsis-10 to determine the event type categories, as it was derived from PubMed Open Access case reports.
-In the first step, event type categories were generated using a language model.
-Since the appropriate categories were not known a priori, we utilized OpenAI’s o3-mini, to analyze all manually annotated clinical events from the sepsis-10 dataset and derive six distinct event type categories.
-The model was prompted to examine the extracted events and construct a comprehensive yet non-overlapping set of categories that could effectively organize clinical events within the dataset.
-The full prompt used for this categorization process is as follows:
-You are a physician and a machine learning scientist specializing in clinical event extraction.
-Your goal is to categorize extracted clinical events into exactly six distinct and mutually exclusive event categories to ensure clarity and consistency.
-Constraints: Each clinical event must belong to exactly one category.
-The categories must be coarse enough to prevent ambiguity.
-The sixth category is reserved for “Other or Unknown.” The first five categories should comprehensively cover the key types of clinical events.
-The categories must be mutually exclusive, meaning an event should unambiguously fit into only one category.
-Task: Identify six well-defined, non-overlapping event categories that best classify the extracted clinical events.
-Format the output as a comma-separated table with two columns: Event Type Category: The name of the category.
-Category Number: An integer from 0 to 5, with “Other or Unknown” always assigned to 5.
-[Category Name], 0
-[Category Name], 1
-[Category Name], 2
-[Category Name], 3
-[Category Name], 4
-Other or Unknown, 5
-Now, based on the extracted clinical events below, generate six non-overlapping event categories and format the output as specified:
-[…]
-The resulting six event type categories, along with their assigned numeric labels, were:
-  • Patient Background and Medical History (0) – Events describing patient demographics, prior medical conditions, risk factors, and relevant history before the clinical presentation.
-  • Clinical Presentation and Examination Findings (1) – Events capturing the patient’s initial symptoms, subjective complaints, and observable clinical signs from physical examination.
-  • Diagnostic Testing and Results (2) – Events related to diagnostic investigations, including laboratory tests, imaging studies, and other diagnostic assessments, along with their results.
-  • Clinical Management and Interventions (3) – Events detailing treatments and medical interventions, including pharmacological therapies, surgical procedures, and other clinical management strategies.
-  • Clinical Course, Outcomes, and Follow-up (4) – Events describing the progression of the condition, response to treatment, complications, recovery, prognosis, and post-treatment follow-up.
-  • Other or Unknown (5) – Events that do not clearly fit into the defined categories or contain insufficient information to be classified.
-Event Type Category Assignment.
-Once these categories were established, they were used to systematically classify each identified clinical event from both sepsis-10 and i2m4.
-The classification was performed separately for clinical events identified through manual annotation and those identified using Llama 3.3.
-To maintain compliance with IRB regulations, all event classifications, including those for i2m4, were executed on a local machine to prevent data transfer to external servers.
-For this classification step, we used Llama 3.3 to assign each extracted clinical event to one of the six predefined event type categories.
-The model was provided with a prompt that explicitly listed all event type categories and required it to assign each event to the most appropriate category.
-The classification was performed for both manually annotated events and events identified by Llama 3.3 across the sepsis-10 and i2m4 datasets.
-The exact prompt used for this classification task is as follows (the green text includes the few-shot examples provided to the model):
-You are a medical professional. You are tasked with categorizing clinical events extracted from case reports.
-Assign the following clinical event to one of these categories.
-The categories are:
-Patient Background and Medical History: 0,
-Clinical Presentation and Examination Findings: 1,
-Diagnostic Testing and Results: 2,
-Clinical Management and Interventions: 3,
-Clinical Course, Outcomes, and Follow‐up: 4,
-Other or Unknown: 5
-The categories are defined as follows:
-Patient Background and Medical History: (Includes demographic details, prior medical diagnoses, past surgical histories, medication use as part of chronic history, and other baseline background information.),
-Clinical Presentation and Examination Findings: (Includes symptoms at presentation, physical exam findings—including vital signs, neurological scores that reflect exam observations—and other immediate clinical observations.),
-Diagnostic Testing and Results: (Includes all imaging studies, laboratory tests, diagnostic procedures and their reported findings, and formal diagnostic conclusions reached via workup.),
-Clinical Management and Interventions: (Includes all treatments, procedures, medications administered acutely, operations, supportive care measures, and decisions/interventions intended to alter the patient’s condition.),
-Clinical Course, Outcomes, and Follow‐up: (Includes statements about change in clinical status, response to treatment, complications, transitions in care, recovery, discharge, and long‐term outcomes.),
-Other or Unknown: (For events that do not clearly fit into any of the above five categories.)
-For example, here is a list of clinical event text and the corresponding category:
-Examples:
-“60-year-old female” → 0
-“history of atrial fibrillation” → 0
-“weighed 95 kg” → 0
-“Impaired consciousness” → 1
-“high-grade fever” → 1
-“persistently high-temperature spikes” → 1
-“Head CT” → 2
-“no hepatitis A” → 2
-“stage IV lymphoma” → 2
-“successfully treated with fidaxomicin” → 3
-“shifted to cefepime” → 3
-“intravenous immunoglobulins for 5 days” → 3
-“follow-up evaluations were recommended” → 4
-“transferred to a geriatric medicine unit” → 4
-“Discharge” → 4
-“he” → 5
-“confined, 5
-“other symptoms” → 5
-[…]
-Respond with only the corresponding integer (0-5) from the list above.
-You have to pick only one category for each event.
-If there is no clear category, choose the category 5 that corresponds to “Other or Unknown” category.
-If there is more than one category, choose the category that you think is most relevant one.
-Do NOT include any extra text in your response.
-Do NOT show your thought process.
-Only provide the integer corresponding to the category and nothing else.
-Format your response as:
-Response: <integer >
-[…]
-Event Type Category Alignment and Match Rate Analysis.
-To assess the consistency of event type categorization between the two annotators (manual annotator vs.
-Llama 3.3), we computed the event type category alignment rate for each dataset separately.
-This metric quantifies how often the manually annotated events and the Llama 3.3-extracted events—when determined to be aligned—are also assigned to the same event type category.
-For each case report in the dataset, clinical events extracted by the manual annotator were aligned with those extracted by Llama 3.3 using an iterative matching approach based on the cosine similarity of their sentence-transformer embeddings.
-Aligned event pairs, which consist of one event from the manual annotator and one from Llama 3.3, were then categorized independently by Llama 3.3 into one of six predefined event type categories of the previous section.
-The event type category alignment rate was computed by determining, for each aligned event pair, whether both events were assigned to the same category.
-If both events in an aligned pair received the same category label, the event type alignment was considered successful.
-For each dataset, the event type alignment rate was first calculated per case report by dividing the number of successfully aligned event type pairs by the total number of aligned pairs in that case report.
-These per-case alignment rates were then averaged across all case reports in the dataset to obtain the final alignment rate.
-Table E.3 reports the mean alignment rate along with the standard deviation for each dataset.
-The event type category alignment rate was higher for the i2m4 dataset ($0.71\pm 0.05$) compared to the sepsis-10 dataset ($0.56\pm 0.05$).
-We hypothesize that this discrepancy is due to the absence of the “Other or Unknown” category in the sepsis-10 dataset, which likely introduced ambiguity in category assignments.
-Specifically, since no events in sepsis-10 were assigned to the “Other or Unknown” category, the model may have redistributed events that would have otherwise fallen into this category among the other five categories.
-This misalignment in categorization likely contributed to the lower agreement rate between the two annotators in the sepsis-10 dataset.
-Dataset Event Type Alignment
-i2m4 $0.71\pm 0.05$
-sepsis-10 $0.56\pm 0.05$
-Table E.3: Mean event type category alignment rate between the manual annotator and Llama 3.3 for each dataset.
-The alignment rate is computed as the proportion of aligned event pairs that were assigned to the same event type category.
-The reported values represent the mean alignment rate across case reports, with standard deviation indicated.
-To further investigate the impact of event type categorization on event match rates, we analyzed the match rate separately for each of the six predefined event categories within both datasets.
-Figure E.4 displays the mean match rate for each event category, with error bars representing the standard deviation across case reports.
-The dashed horizontal lines indicate the overall mean match rate for all event types within each dataset (as seen in Table 2), providing a point of reference for comparison.
-From the figure, it is evident that the event match rates exhibit variability across categories but do not show significant deviations that would suggest one category consistently performs better or worse than the others.
-For instance, in both datasets, categories such as “Diagnostic Testing and Results” and “Clinical Course, Outcomes, and Follow up” tend to achieve higher match rates compared to others, though the differences are not substantial enough to be statistically significant.
-Similarly, the match rates for “Clinical Management and Interventions” and “Patient Background and Medical History” fall closer to the overall dataset mean, suggesting these categories are neither particularly challenging nor exceptionally easy for alignment.
-A notable observation is the “Other or Unknown” category, which is entirely absent in the sepsis-10 dataset.
-This absence introduces a unique challenge for categorization, as events that might have otherwise been assigned to this category are redistributed among the remaining categories.
-The i2m4 dataset, on the other hand, includes this category, and has a moderate event match rate for the event types in this category.
-This discrepancy could partially suggest that the style, structure, or format of the sepsis-10 dataset differs from that of i2m4.
-The absence of the “Other or Unknown” category in sepsis-10 could indicate a more standardized or narrowly focused structure, where fewer ambiguous or unclassifiable events are present, whereas i2m4 may include a broader range of events requiring such a fallback category.
-This difference in dataset characteristics could contribute to the observed variability in event match rates, as the annotation and alignment processes are inherently influenced by the dataset’s structure and complexity.
-Figure E.4: Mean event match rate across predefined event type categories for the i2m4 and sepsis-10 datasets.
-Bars represent the mean match rate for each event type category, calculated as the proportion of aligned event pairs (between the manual annotator and Llama 3.3) with cosine similarity less than the matching threshold.
-Error bars indicate the standard deviation of match rates across case reports.
-The dashed lines represent the overall mean match rates for each dataset across all categories (From Table 2: sepsis-10: 0.753 and i2m4: 0.404) The “Other or Unknown” category is absent in sepsis-10, resulting in no match rate for this category in that dataset.
-Overall, the analysis suggests that event match rates are relatively stable across event categories within each dataset.
-The differences in overall match rates between i2m4 and sepsis-10 are therefore more likely attributable to dataset-specific characteristics and under-identification bias that was explained in the previous section.
-Future work could explore more granular metrics, such as examining the semantic content of events within each category, to better understand the nuances of these event match rate discrepancies.